### v1

# Integration of TinyML and LargeML: A Survey of 6G and Beyond

Abstract
The transition from 5G networks to 6G highlights a significant demand for machine learning (ML). Deep learning models, in particular, have seen wide application in mobile networking and communications to support advanced services in emerging wireless environments, such as smart healthcare, smart grids, autonomous vehicles, aerial platforms, digital twins, and the metaverse. The rapid expansion of Internet-of-Things (IoT) devices, many with limited computational capabilities, has accelerated the development of tiny machine learning (TinyML) and resource-efficient ML approaches for cost-effective services. However, the deployment of large-scale machine learning (LargeML) solutions require major computing resources and complex management strategies to support extensive IoT services and ML-generated content applications. Consequently, the integration of TinyML and LargeML is projected as a promising approach for future seamless connectivity and efficient resource management.
Although the integration of TinyML and LargeML shows abundant potential, several challenges persist, including performance optimization, practical deployment strategies, effective resource management, and security considerations. In this survey, we review and analyze the latest research aimed at enabling the integration of TinyML and LargeML models for the realization of smart services and applications in future 6G networks and beyond. The paper concludes by outlining critical challenges and identifying future research directions for the holistic integration of TinyML and LargeML in next-generation wireless networks.
Index Terms:
6G, artificial intelligence (AI), deep learning (DL), distributed learning, federated learning (FL), Internet-of-Things (IoT), large machine learning (LargeML), machine learning (ML), tiny machine learning (TinyML).
I Introduction
The evolution of communication technologies has been characterized by extraordinary advancements every decade. From the first-generation (1G) analog systems, which introduced mobile voice communications, to fifth-generation (5G) networks currently transforming industries with ultra-fast speeds and low latency, each generation has delivered major transformative impacts [1]. Looking ahead, sixth-generation (6G) networks are expected to push technological boundaries further, offering unprecedented capabilities and driving global innovation. The core features of 6G will clearly differentiate it from previous generations (Fig. 1), positioning the paradigm as a transformative technology for diverse applications. These features address the escalating demands for higher data rates, superior reliability, efficient spectrum management, and superintelligence capabilities. Furthermore, 6G aims to integrate advanced artificial intelligence (AI) and ML technologies seamlessly, expanding the possibilities of wireless networks and significantly improving the delivery of high-end applications and services to users.
Figure 1: Transition from 5G to 6G and beyond (B6G): From target throughputs to use cases, core technologies, and emerging applications.
The integration of AI and ML into 6G will enable new levels of automation and intelligence, unlocking numerous new applications and high-end services [2]. AI already operates on multiple network layers to improve dynamic resource management, predict maintenance needs, and personalize user experiences. Advances in ML, particularly deep learning (DL) architectures that execute computer vision and natural language processing (NLP) tasks [3], are expected to support high-fidelity holographic communications. This capability will allow users to interact in immersive three-dimensional environments and could revolutionize remote collaboration. In a hyper-connected future, 6G will unlock the full potential of Internet-of-Things (IoT) and mobile big data to drive innovation and lead the world into smarter, deeper connectivity.
I-A Context and Motivation
Table I: Comparison of TinyML and LargeML Features.
Aspects TinyML LargeML
Definition ML on resource-constrained devices Large-scale AI models with vast numbers of parameters and large data processing capacity
Computational requirements Limited power, memory, and processing capability Substantial power, memory, and computational resources
Deployment environment Smart sensors, edge devices, and microcontrollers Cloud servers and specialized hardware (e.g., GPUs, TPUs)
Latency Ultra-low latency for real-time processing Higher latency, acceptable for batch processing
Power consumption Very low, suitable for battery-powered devices High, requiring a robust power supply
Data processing Local (on-device) processing Centralized (large-scale) data processing
Typical applications Real-time analytics and sensor data processing in IoT environments (e.g., smart homes, industrial monitoring) Complex data analysis and decision-making (e.g., NLP, image recognition, autonomous systems)
Advantages Low latency, reduced data transmission, enhanced privacy High accuracy, with the ability to process and analyze large datasets
Challenges Constrained by hardware limitations and task complexity Requires extensive computational infrastructure and incurs high costs
Complementary roles Supports real-time, lightweight, and edge-based tasks Enables deep, comprehensive analysis and cloud-based decision-making
Examples Keyword spotting, anomaly detection in sensor data GPT-4, BERT, large-scale image classifiers
In the last decade, AI has converged around two contrasting approaches: TinyML and Large ML (LargeML). TinyML focuses on creating efficient models for resource-constrained edge devices, allowing real-time analytics close to data sources (i.e., bringing analytics closer to data sources) [4]. In contrast, LargeML leverages powerful cloud servers to handle complex, data-intensive tasks, exemplified by models like GPT-4 and BERT, which excel in applications such as sentiment analysis and creative text generation [5] (e.g., Large Language Models (LLMs), generative AI (GenAI), and Agentic AI).
However, LargeML’s reliance on significant computational resources limits its real-time processing abilities on constrained devices and raises privacy concerns. TinyML mitigates these issues by pre-processing data at the edge, while LargeML can deploy compressed models on edge devices through knowledge distillation [6]. This collaborative learning strategy permits more flexible deployment and operation of IoT applications and services. TinyML benefits from the sophisticated learning capabilities of LargeML to improve accuracy, whereas LargeML gains from the real-time processing potential and privacy-preserving features of TinyML. Table I presents a concise comparison of TinyML and LargeML.
The integration of TinyML and LargeML addresses these individual limitations by combining their strengths, enhancing AI capabilities in 6G and future networks. For instance, wearable health monitors can use TinyML for real-time vital sign analysis [7], while LargeML can detect medical risks using historical data. Similarly, in virtual environments, TinyML can process wearable Electroencephalography (EEG) data to gauge emotional states [8], with LargeML translating this into actions within the metaverse [9].
Despite these advantages, the integration faces challenges, including increased security risks from numerous connected devices and ethical concerns regarding AI explainability and data privacy. In summary, leveraging the unique strengths of TinyML and LargeML can lead to seamless device-to-server interactions and drive advancements in data-driven innovation and intelligent living environments.
I-B State-of-the-Art and Contributions
The AI ecosystem is evolving rapidly, driven in part by two distinct yet complementary paradigms: TinyML and LargeML. TinyML enables resource-limited edge devices to perform on-device ML tasks with local data, whereas LargeML employs powerful server infrastructures to train complex models on vast datasets. As the IoT continues to evolve and the advent of 6G approaches, the collaborative potential of these two different learning strategies presents a viable path on which to increase the performance of IoT applications and 6G network operations, especially under challenging conditions. This section reviews the recent survey literature on TinyML and LargeML systems, highlighting their contributions and limitations in the context of 6G and its envisioned applications.
Over the past decade, numerous survey works have examined the role of TinyML in edge computing. Key contributions [10, 11, 12, 13, 14, 15] provide comprehensive overviews of TinyML, covering its fundamentals, development tools, applications, and future directions. These surveys highlight both the transformative potential and widespread use of TinyML in domains such as industrial IoT, smart healthcare, autonomous driving, environmental monitoring, public safety, human-machine interaction, agriculture, and emergency response. A significant contribution noted in [10] is the concept of hardware-software cooperative design for efficient TinyML implementation. This survey highlights the necessity of optimizing both hardware and software components concurrently for efficient TinyML execution on resource-constrained devices. Coordinated design enhances model efficiency while minimizing power consumption and computational overhead. Similarly, the survey in [12] offers a systematic overview of existing TinyML development tools, including hardware platforms, software frameworks, and supporting libraries. These resources assist AI developers and specialists in designing and deploying TinyML models at the network edge. Other surveys [11, 14, 15] explore a range of TinyML applications across domains such as anomaly detection, predictive maintenance, and intelligent sensor networks. For example, anomaly detection allows TinyML models to continuously monitor sensor data and identify deviations from normal operating behavior, supporting timely maintenance and minimizing unplanned downtime. In predictive maintenance, TinyML models forecast potential equipment failures, improving resource allocation and operational efficiency through proactive maintenance strategies. Intelligent sensor networks use TinyML to process data locally at the edge, enabling quicker decision-making and near real-time responsiveness. These applications illustrate the broad potential of TinyMl to transform sectors such as industrial automation, environmental monitoring, and smart infrastructure.
Despite these advances, the existing survey literature on TinyML reveals some limitations when considered in the context of 6G integration. Although these surveys examine TinyML applications within general IoT frameworks, they overlook the specific challenges and opportunities presented by integrating TinyML with LargeML in 6G networks [11]. Some surveys [14, 15] provide limited clarification on fine-tuning TinyML models to the specific resource constraints of 6G environments. For example, 6G may require the development of lightweight communication protocols tailored to TinyML models for efficient data exchange with server-side LargeML systems. These surveys also fall short of exploring optimization strategies for minimizing power consumption in TinyML models while preserving the accuracy required for effective collaboration with LargeML [15].
Table II: Summary of related surveys on TinyML and LargeML.
Refs. Review Topics Key Contributions Limitations
TinyML LargeML
[10] Highlights hardware-software co-design for optimizing TinyML frameworks on resource-limited devices and reviews ML algorithms optimized for low-power deployment. Concentrates on TinyML and 5G, without discussion of integration with LargeML in 6G systems.
[11] Introduces a classification system for TinyML applications based on an extensive literature review, exploring benefits and use cases. Does not address TinyML–LargeML integration or the specific challenges of 6G network integration.
[12] Surveys TinyML frameworks, development tools, enablers, and applications and outlines current research challenges. Lacks discussion of TinyML–LargeML integration in the context of 6G networks.
[13] Proposes a taxonomy for reformable TinyML solutions and evaluates the suitability of TinyML layers for reformability. Does not explore TinyML–LargeML integration in the context of 6G networks.
[14] Reviews neural network (NN) compression techniques and categorizes TinyML hardware and software resources. Examines the foundational aspects of TinyML, with limited discussion on integration with LargeML or implications for 6G environments.
[15] Reviews TinyML research, workflows, algorithms, and development tools, with an emphasis on recent ML advancements. Lacks analysis of integration with LargeML in the context of 6G networks.
[16] Reviews transformer architectures in LLMs, discussing training methods, applications, societal impact, and deployment challenges. Does not address 6G-related challenges or TinyML–LargeML integration for enhancing 6G capabilities.
[17] Examines LLM evaluation methods, benchmarks, and performance across domains. Omits discussion of TinyML–LargeML integration and analysis of challenges in enhancing 6G applications.
[18] Analyzes fine-tuning algorithms for LargeML, detailing performance metrics and deployment scenarios. Does not investigate the integration of fine-tuning algorithms with TinyML for 6G applications.
[19] Proposes a taxonomy of explainability techniques for LLMs, with evaluation metrics and implementation challenges. Concentrates on LLM explainability without addressing its integration with TinyML in collaborative 6G systems.
[20] Outlines architectures, techniques, and challenges in pre-trained language models (PLMs). Concentrates on natural language processing (NLP) applications of PLMs; lacks discussion of TinyML integration in 6G systems.
This work Offers the first survey of the integration of TinyML and LargeML for 6G systems. Specifically: Introduces fundamentals of TinyML and LargeML. Discusses motivations and requirements for integrating TinyML and LargeML in 6G. Proposes bidirectional integration frameworks for enhancing future network performance and capabilities. Explores potential applications of integrated TinyML–LargeML systems. Identifies challenges and future research directions for TinyML–LargeML integration.
In contrast, many studies on LargeML outline distinct contributions and limitations. Surveys such as [16, 17, 18, 19, 20], explore the architectures, applications, evaluation methods, and research challenges associated with LargeML. These works offer a clear view of the current state of the LargeML field. For example, the work in [16] presents a detailed analysis of transformer-based architectures, from foundational models to advanced, with a focus on LLMs. It traces the development of of LLMs, discussing training methods, application domains, and real-world deployments. Similarly, the survey in [20] explores the impact of large pre-trained language models (PLMs) on a range of NLP tasks. These surveys highlight the strong performance of LargeML in complex tasks such as text generation, machine translation, and sentiment analysis. Several reviews [17, 18, 19] examine key challenges associated with LargeML. These include explainability, where understanding model decision-making processes remains critical for real-world deployment. Notably, [18] emphasizes the need for parameter-efficient fine-tuning (PEFT) techniques to reduce computational overhead, facilitating deployment in resource-constrained environments. Addressing these challenges is essential for the responsible and effective use of LargeML across diverse applications. However, as with the surveys on TinyML, current LargeML reviews exhibit limitations when considered in the context of 6G applications. Although the computational demands of LargeML are acknowledged, the specific challenges of deployment and and coordination with LargeML within 6G networks are not examined. Given 6G’s focus on real-time processing and edge intelligence, it becomes necessary to investigate how LargeML models can be compressed or adapted for efficient operation at the network edge, potentially in conjunction with TinyML systems. Furthermore, the surveys overlook the impact of 6G-specific features such as novel communication protocols and strict latency constraints on the training and deployment of LargeML models.
The preceding discussion highlights a significant gap in comprehensive surveys that examine the integration of TinyML and LargeML in the context of 6G and beyond. Although recent progress in LargeML integration [21, 22] is critical for developing efficient and intelligent 6G networks, these advancements have not been thoroughly examined. Existing studies focus on isolated aspects of either TinyML or LargeML [23], failing to provide a cohesive analysis of their combined potential to address emerging network challenges. The present survey aims to fill this gap by presenting a comprehensive survey on the integration of TinyML and LargeML for 6G and future network paradigms. A key contribution is a detailed examination of the motivations and requirements behind such integration. We begin by reviewing the foundations and recent advancements in both TinyML and LargeML, highlighting their respective strengths. We then propose efficient bidirectional integration strategies and asses their relevance to future network architectures. In addition, we examine the real-world applications of integrated TinyML–LargeML systems in areas such as smart healthcare, autonomous vehicles, and industrial IoT. These case studies illustrate the potential to improve communication efficiency, optimize resource use, and scale systems systems effectively. Finally, we identify major unresolved challenges in the literature and propose research directions to guide future work. The features and contributions of this work are summarized below:
  • Background: This work first provides the background of TinyML and LargeML, accordingly highlighting their individual advancements and potential cooperation.
  • Integration requirements: We outline the motivations and requirements for integrating TinyML and LargeML to emphasize their importance for 6G networks.
  • Efficient solutions: We propose and discuss efficient bidirectional integration solutions, aiming to enhance the performance and capabilities of future networks.
  • Applications of integrated systems: We conduct an extensive review of the applications of integrated TinyML and LargeML systems across various domains to demonstrate their practical benefits and potential impacts.
  • Discussion of challenges and future research: From our extensive review, we identify critical challenges and propose several potential directions for future research, including resource allocation, communication efficiency, interplay with LargeML, and standardizations.
I-C Paper Organization
The remainder of the paper is structured as follows. Section II provides an overview of TinyML and LargeML, outlining their key concepts and recent developments Section III examines the motivations and requirements for integrating TinyML and LargeML in the context of 6G networks. Section IV presents a comprehensive review of efficient, bi-directional integration approaches for 6G and beyond, such as transfer learning (TL), federated transfer learning (FTL), split learning (SL), and federated split learning (FSL). Section V explores the applications of integrated TinyML and LargeML systems in domains such as data privacy and network security, network management, zero-touch networks, and the brain-level metaverse. In Section VI, we summarize insights from earlier sections and discuss existing challenges and potential directions for future research before ending with Section VII.
II Overview of TinyML and LargeML
Figure 2: Overview of TinyML and LargeML: life cycle, design, operation, and performance evaluation.
II-A TinyML
TinyML refers to a lightweight, resource-constrained class of ML designed for environments with limited memory and processing capabilities [4]. Despite its compact form, TinyML performs effectively in a wide range of communication tasks; Table III details its characteristics.
Table III: TinyML classifications, hardware devices, environmental deployment, and software tools.
Classification Hardware Device [15] Environmental Deployment [11] Software Tools [13]
Model size: (1) Compression techniques (pruning, quantization, knowledge distillation), (2) Memory and storage design (kilobytes to megabytes), (3) Operational modes (centralized, distributed, decentralized). Architecture: (1) Convolutional neural network (CNN) for image processing, (2) Recurrent neural network (RNN) for sequential data, (3) Fully connected networks for certain tasks, (4) Transformer [24] for data generation, context awareness, translation, and summarization. Training data: (1) Quality assurance and data labeling, (2) Data augmentation techniques (rotation, scaling, flipping), (3) Fine-tuning pre-trained models on large datasets. Algorithms: Ten types [11]: supervised, weakly supervised, unsupervised, self-supervised, meta-learning, continual learning, deep reinforcement learning (DRL), TL, SL, and federated learning (FL). Central processing unit (CPU): Arm Cortex M family offers ultra-low-power operation but limited support for parallel processing. Graphics processing unit (GPU): Supports parallel computation and optimized sequential processing of large datasets (e.g. NVIDIA Jetson family, AMD, Intel, Arm), but is not cost-effective. Field-programmable gate array (FPGA): Enables model customization and hardware-level acceleration for complex operations and large dataset access, but lacks broad library support. Tensor processing unit (TPU): Accelerates ML workloads, particularly those involving NNs, with two representations of Edge TPUs and Cloud TPUs. Atmospheric: Low-power sensors to monitor air quality in urban, forest, and industrial areas, detecting pollutants by measuring particulate matter, carbon dioxide levels, and others. Hydrosphere: Sensors in aquatic environments to detect pollutants, measure pH, and monitor temperature changes. Biosphere: Device used for wildlife conservation, monitoring biological activity, or classifying sounds (e.g., animal habitats and anti-poaching surveillance). Lithosphere: Applied in agriculture for environmental monitoring (e.g., soil moisture, nutrient levels, pests) or earthquake detection. Human-related: Wearables for health monitoring (e.g., blood pressure) and assistive technologies for humans (e.g., mobility). Mobile device: TensorFlow Lite for model optimization and PyTorch for programming. Embedded devices: TensorFlow Lite for model optimization, PyTorch for programming, and Qeexo AutoML for end-to-end customization. Edge devices: PyTorch for programming, NanoEdgeAIStudio for low/no-code solution, and Imagimob/Edge Impulse for end-to-end customization. Microcontroller: Arm NN for network architecture optimization, MinUn for bandwidth/memory management, STM32CubeA and uTensor for pre-trained model conversion, and NXP eIQ as a development environment.
According to [25], the TinyML life cycle consists of three stages:
  • Stage 1: Data collection from peripheral devices, such as sensors or production environments (real or simulated).
  • Stage 2: Model deployment using quantization and compression techniques to address hardware heterogeneity and resource constraints.
  • Stage 3: Model operationalization by converting the trained model into a generic, interpretable form suitable for target devices.
TinyML relies on application-specific data formats, prioritizing model customization for limited device capacity and efficient data transmission. The next sections detail these design strategies and operational considerations (Fig. 2 for short overview).
II-A1 TinyML Design
Although no standard guidelines for TinyML design currently exist, several best practices are commonly adopted to guide development.
Quantization: Quantization reduces the model’s memory footprint and accelerates computation by representing parameters with lower precision numerical formats (e.g., converting 32-bit floating point values into 8-bit integers) [26]. Quantization can be implemented during training (quantization-aware training) or after training (post-training quantization). Variants include uniform and non-uniform quantization for optimizing accuracy, symmetric or asymmetric schemes to reduce quantization error, and Bayesian inference for minimizing bit encoding requirements.
Pruning: Pruning reduces NN size by removing unimportant parameters, such as weights, neurons, or layers, that contribute little to model performance [27]. Pruned models typically require fine-tuning to restore accuracy. Common strategies include threshold-based pruning, in which weights below a specified magnitude are discarded. Depending on the differences in the model weight matrix, pruning is classified as unstructured (removal of individual weights) or structured (removal of neurons, filters, rows, or columns), guided by redundancy or regularization metrics.
Low-rank matrix decomposition: This technique applies linear algebra methods, such as singular value decomposition or its extensions to tensors, to approximate large weight matrices with lower-rank components. The resulting factorized forms contain fewer parameters, reduce computational complexity, and simplify matrix operations [28]. Despite its benefits, this technique can be computationally intensive and more difficult to implement, as the decomposed network requires additional fine-tuning to recover accuracy loss, similar to model pruning.
Weight sharing: Weight sharing reduces the number of unique parameters for storage and computation by reusing weights across different parts of the model [29]. For example, shared weights can be arranged by row or column within a weight matrix. However, designing effective sharing configurations manually is difficult and can lead to unpredictable performance, despite the inherent redundancy in DL models.
Algorithm-architecture co-design: Integrating algorithm development with specialized hardware accelerators, such as TPUs or custom-designed chips (e.g., Cortex-M7, Cortex-M4, Cortex-M0+, and eDMPv1), can significantly improve the inference efficiency in TinyML models[30]. This joint optimization of algorithms and hardware also enhances computational performance and energy efficiency.
Huffman coding: Huffman coding assigns variable-length codes to symbols based on their frequencies, with more frequent symbols receiving shorter codes [31]. When applied to model weights and parameters, this technique can significantly reduce overall model size through efficient encoding.
Hyperparameter optimization: This refers to the process of selecting the optimal configuration of model parameters, such as learning rate, batch size, activation functions, number of neurons, and network depth [32]. Common optimization methods include grid search, random search, Bayesian optimization, Hyperband (which dynamically allocates resources to promising configurations while terminating underperforming configurations early), genetic algorithms, tree-structured Parzen estimators, and hyperdimensional computing. These techniques aim to improve model performance by minimizing memory and computational overhead while ensuring user-defined accuracy.
Knowledge distillation: Knowledge distillation involves transferring learned representations from a larger model (the teacher) to a smaller model (the student) to produce a compact model with comparable performance. Three common types of distillation techniques are (1) response-based, where the student is trained to mimic the teacher’s final output layer, (2) feature-based, where the student replicates intermediate feature representations from the teacher, and (3) relation-based, where the student learns the structural relationships between different components of the teacher model [33].
II-A2 TinyML Operation
From an operational standpoint, TinyML can be categorized into three main levels: firmware updates, model operation, and communication protocols [34].
Firmware updates: This involves replacing the entire software stack on a device, including the TinyML model and underlying system components. Although comprehensive, this process is resource-intensive and time-consuming. Firmware updates are typically delivered over-the-air (OTA) using lightweight machine-to-machine (M2M) protocols, which support both push and pull modes for firmware transfer [35]. Tools such as RIOT-ML facilitate secure OTA updates, enabling remote, unattended deployment.
Model operations: LargeML deployment involves multiple strategies to optimize model updates and execution on resource-constrained devices [34]. Full model updates replace the entire ML model via OTA delivery, ideal for significant architecture or parameter changes, ensuring devices run the latest version. Partial model updates target specific layers or parameters, minimizing network traffic and power use, making them suitable for fine-tuning on constrained devices [36]. Model transpilers convert models between frameworks, like TensorFlow to TensorFlow Lite Micro, enhancing compatibility and resource efficiency across diverse hardware [37]. Model compilers like uTensor and TinyTS translate models into optimized binary code for microcontrollers, using quantization and pruning to reduce size and boost inference efficiency, fitting within tight memory and computational limits.
Communication protocols: TinyML systems leverage specialized communication protocols to optimize data transfer under power, bandwidth, and range constraints in IoT ecosystems [15]. For intra-device communication, protocols like inter-integrated circuit and serial peripheral interface enable efficient data exchange between microcontrollers and peripherals during local collection. Device-to-cloud communication often employs message queuing telemetry transport, a lightweight protocol ideal for low-bandwidth, high-latency environments. Short-range wireless needs are met by low-power wide-area network (LoWPAN) protocols such as Bluetooth, Zigbee, or 6LoWPAN for wearables and smart home sensors, while Wi-Fi supports higher data volumes in smart cameras or environmental monitors. For long-range, low-power communication, long-range wide area network excels in remote applications like agriculture and smart cities, offering kilometer-scale coverage with minimal energy use. Sigfox, designed for ultra-narrowband IoT and M2M tasks, transmits small, infrequent data packets on unlicensed bands, ensuring low-cost, secure deployments. Similarly, Narrowband IoT, a 3GPP standard, supports low-data-rate, energy-efficient communication in licensed/unlicensed spectra, enabling scalable, reliable connectivity for massive IoT deployments in challenging environments like underground or indoor settings.
II-A3 Performance Evaluation
Given the diversity of application contexts and strict resource limitations, performance evaluation of TinyML models requires the consideration of both device constraints and learning model characteristics.
Memory Footprint: TFLite Micro targets microcontrollers with limited ROM and RAM, using quantization and pruning to reduce model size with minimal accuracy loss [38]. Evaluating ROM and RAM consumption is crucial for assessing TinyML deployments on IoT devices.
On/Off characteristics: TinyML components (sensors, protocols, inference engines) switch between sleep and active power states to save energy [37]. Optimizing the duration and frequency of these transitions minimizes active time, enhancing energy efficiency while preserving functionality.
Complexity profiling: This refers to the analysis of the computational demands of TinyML models. Floating-point operations are more computationally intensive than quantized int8 operations [15], which reduce weight and activation precision to shrink model size and accelerate inference. Complexity profiling helps quantify the trade-offs between model size, accuracy, and inference performance [4].
Energy profiling: Energy profiling measures CPU cycles, latency, and total energy consumption during inference [39]. Floating-point models typically consume more energy than quantized int8 models due to the greater computational overhead. Thus, it helps identify energy-efficient TinyML models.
Model Evaluation Metrics: Several metrics are used to evaluate TinyML model performance: Accuracy, mean absolute error (MAE), root mean square error (RMSE), R-Squared, precision, recall, F1-score, and cross-entropy loss.
Deployment robustness: Deployment robustness ensures TinyML model reliability across environmental conditions. Two evaluation approaches are used [34]: per-model evaluation assesses performance in controlled pre-deployment tests, while per-operator evaluation monitors adaptability in real-world settings. Robustness testing involves diverse datasets, hardware, and environments to ensure consistent performance.
II-B LargeML
Integrating LargeML into wireless networks is a key trend for enabling autonomous operations, using cognitive modules for resource allocation, channel estimation, and mobility management [5]. The LargeML life cycle consists of three stages:
  • Stage 1: Collect and pre-process datasets from multiple domains to construct a high-quality training corpus.
  • Stage 2: Design model architectures by selecting appropriate layer types and arrangements, configuring parameters, and determining suitable training algorithms.
  • Stage 3: Train the model in two phases: (1) Pre-train on broad datasets to build generalized representations; (2) Fine-tune on task-specific data for applications such as question answering, summarization, and code generation.
  • Stage 4: Evaluate performance using established benchmarks. Refine model architecture, hyperparameters, and datasets based on results.
  • Stage 5: Apply model compression techniques to reduce the memory footprint and inference latency without sacrificing performance.
  • Stage 6: Test and optimize the model in controlled environments to maximize inference efficiency without degrading output quality.
  • Stage 7: Deploy in production settings. Continuously monitor, fine-tune, and adapt the model to accommodate evolving data patterns and user requirements.
  • Stage 8: Mitigate risks and enhance transparency by filtering outputs through techniques such as adversarial training, fairness constraints, and curated datasets. These steps address bias, privacy, security, and misuse concerns.
II-B1 LargeML Design
The design process for LargeML covers six key areas: data processing, learning architecture, network configuration, pre-training tasks, fine-tuning strategies, and prompt engineering. Each area is outlined below.
Data Processing
Training LargeML relies on extensive IoT filme data, where data quality significantly impacts prediction and decision accuracy. Effective data processing involves key actions to ensure high-quality outcomes.
Collection and labeling: Data is gathered from heterogeneous sources (i.e., real-time sensors, network logs, user feedback, and environmental inputs). These data points are annotated with application-relevant features (e.g., signal strength, noise levels, and user mobility patterns [40]). Consistent and accurate labeling is critical to prevent bias in state or event label representations [41]. Well-structured datasets enable the model to generalize across a range of operational scenarios.
Cleaning, transformation, and augmentation: This phase prepares the raw data for modeling. Cleaning involves removing duplicates, correcting errors, handling missing values, and filtering noise. Transformation includes normalization, encoding categorical variables (e.g., one-hot or label encoding), and feature engineering. These transformations convert raw data into model-ready formats for analysis, such as token-based input (for textual and code sequences), tree or graph-based structures (to capture syntactic or hierarchical relationships), pixel-based representations (for visual encoding of data), and hybrid formats that integrate multiple representations [17]. Augmentation techniques, including rotation, flipping, or noise injection, generate additional samples to enhance model robustness. The overall process ensures a clean, consistent, and well-structured dataset for downstream tasks.
Featuring and splitting: This phase involves feature selection, feature extraction, and data partitioning [42]. Feature selection identifies the variables with the greatest impact on model performance, a process often referred to as sensitivity analysis. Feature extraction derives new variables, for example the rate of change in wireless signal strength, to reveal deeper patterns within the data. After the features are defined, the dataset is partitioned into training, validation, and testing sets to evaluate the model’s performance and prevent bias. Cross-validation, involving training the model on different subsets of the data, is used to assess robustness and generalizability.
Learning Architecture
Future 6G networks aim to support integrated sensing and communication (ISAC), task-oriented communications, and digital twins, requiring LargeML models for efficient real-time data processing [5]. Selecting suitable learning architectures is challenging due to application-specific constraints. Transformers, leveraging attention mechanisms, outperform RNNs in sequence data processing [24]. They use an encoder-decoder structure with stacked sub-layers: encoders create context-aware representations, positional embeddings preserve token order, and decoders apply masked or multi-head self-attention to assess token relevance. Aside from three typical architectures (Encoder only [43], Encoder-Decoder [24], and Decoder only with types of causal decoders [44] and decoders [45]), below are new variants of transformer architectures.
Mixture of experts (MoE): This technique extends all three architectural types by scaling a subset of NN weights, enabling sparse activation. Models such as Switch Transformer [46] and GLaM [47] demonstrate MoE’s ability to increase model parameter counts without proportional increases in computational cost. Performance improves with more experts or larger total parameter size. However, training can become unstable due to complex routing operations. To address this, techniques such as incorporating high-precision tensors in the routing module, narrowing initialization ranges, or introducing resource-conditional computation alongside traditional data-conditional computation have shown promise [48].
Emergence: Conventional transformer architectures may struggle with quadratic computational complexity with respect to sequence length, resulting in high processing costs for long inputs. This limitation has prompted research into parameterized state space architectures [49], which integrate features of RNNs and CNNs. These models generate outputs recursively in the same manner as RNNs, referencing only the previous state during decoding while avoiding the need to revisit all previous states, which is typical in conventional transformers. These models also encode entire sequences in parallel using convolution operations, exploiting GPU efficiency through techniques such as parallel scan, fast Fourier transform, and chunkwise recurrent processing.
Network Configuration
Network configuration directly impacts model training and performance. Selecting the appropriate operational strategy for the model is a critical early step. The following approaches highlight important considerations for efficient network design [50].
Parallel training: This technique accelerates training through parallelism strategies. Data parallelism synchronizes GPUs using a parameter server/all-reduce gradients. Pipeline parallelism divides the model across devices to overlap the execution of data processing. Model parallelism distributes layers or neurons across devices, with each device processing parts of the model and exchanging intermediate outputs.
Mixed precision: During model training, parameter values typically remain within the range of 16-bit floating-point numbers. However, the product of gradients and the learning rate during updates can exceed this range, leading to underflow and ineffective parameter updates. To prevent this problem, optimizers store an auxiliary set of parameters in single-precision (32-bit). While training uses half-precision (16-bit) for parameters and gradients, the optimizer computes and stores updates in 32-bit before conversion into 16-bit.
Offloading: Optimizer parameters often consume more memory than model parameters, requiring offloading from GPU to CPU. However, offloading to a single CPU reduces the training throughput. This bottleneck is overcome by distributing the load across multiple CPUs linked to the GPU.
Overlapping: Memory operations are generally asynchronous. To improve computational efficiency during the model’s forward propagation, a two-phase operation is applied. Memory requests (parameter retrieval) are issued first, followed by computations (parameter processing), allowing the memory operation to complete (new parameter collection).
Checkpointing: In transformer models, backward propagation uses a checkpoint mechanism to optimize memory by saving only selected inputs (checkpoints) at each major layer (attention and feed-forward) during forward propagation, rather than storing all intermediate results. During backpropagation, gradients are computed by re-executing the forward pass for each layer using these checkpoints, after which both checkpoints and recomputed results are discarded, balancing memory efficiency with computational cost.
Pre-training Tasks
Pre-training establishes the groundwork for the model by equipping it with essential skills and knowledge that require minimal task-specific training, involving extensive data collection and preparation, self-supervised learning to acquire foundational knowledge, and fine-tuning for specific tasks. Before deployment, LargeML undergoes pre-training on extensive corpora to learn universal representations that encode both syntactic and semantic information. The models are then fine-tuned for downstream tasks, facilitating knowledge transfer to a broad range of applications. Below are common pre-training workflows [51].
Single modal: This method trains models on a single data modality (e.g., text, images, audio) to build domain-specific representations. In NLP, for example, pre-training on large text corpora enables models to capture linguistic structures and semantics, excelling in tasks such as sentiment analysis, speech signal classification, and machine translation.
Multiple modal: This method combines different data types (e.g., text, images) to build richer representations. For example, training on both image captions and corresponding images enables models to link visual and textual information, improving performance in tasks such as image captioning, visual question answering, and cross-modal retrieval.
Token prediction: Token prediction tasks train models to predict the next token in a sequence, strengthening contextual interpretation and language processing. Masked language modeling (e.g., BERT) hides random tokens for prediction, while autoregressive modeling (e.g., GPT) predicts the next token based on prior content, supporting coherent and contextually accurate outputs.
Fine-tuning
After pre-training, a model is fine-tuned on a smaller, task-specific dataset to adjusts its parameters for the target application. This process builds on the knowledge acquired during pre-training. In addition to the TinyML strategies discussed in Section II-A1, this section reviews the fine-tuning techniques relevant to LargeML systems [52].
Prompt tuning: Prompt tuning introduces trainable tokens (prompts) to guide the model’s output without modifying its architecture or parameters.
Prefix tuning: This technique prepends a set of trainable tokens (prefixes) to the model’s input and intermediate layers. These prefixes steer the model’s attention and interpretative processes without updating the core parameters, enabling task adaptation with low computational overhead.
Adapter tuning: Adapter tuning inserts small NN modules (adapters) into the model’s architecture. During fine-tuning, only the adapters are trained to capture task-specific information [53], while the original model remains unchanged.
Low-rank adaptation (LoRA): LoRA enhances transformer models by introducing trainable low-rank matrices to selected layers, keeping original weights frozen [54]. Only these lightweight matrices are trained to approximate weight updates, enabling efficient fine-tuning with minimal memory and computational costs.
Knowledge preservation: This technique adapts pre-trained models to new tasks while retaining their valuable pre-existing knowledge. It addresses catastrophic forgetting, using techniques such as causal graph framing to preserve important causal effects from pre-trained data, and PEFT techniques such as LoRA and Half Fine-Tuning. This balance ensures that the model remains effective across diverse tasks without overfitting to new data.
Task orientation: This technique adapts pre-trained models to specific tasks (e.g., dialogue systems) by fine-tuning the model’s parameters on task-specific data. The incorporation of techniques such as adapter layers and prefix tuning reduces the overall parameter count for training, improving efficiency and lowering resource demands.
Prompt Engineering
After fine-tuning, prompt engineering refines how the model interprets and responds to inputs. Prompting techniques vary depending on task requirements and application contexts. Common types include multi-turn instructions, multiple modeling, task planning, tool augmented alignment, and retrieval augmentation. The most widely used prompting techniques in wireless networks are described below [5].
Few-shot prompting: This method enables the model to generalize from a small number of examples. It is particularly useful when large annotated datasets are unavailable, allowing the model to perform new tasks with minimal additional data.
Zero-shot prompting: This method exploits the model’s pre-trained knowledge to perform tasks without prior examples or specific training. The model generates outputs based solely on instructions, drawing from its existing knowledge. In contrast to few-shot prompting, zero-shot prompting provides no examples to guide the model’s output.
In-context learning: This method refers to a model’s ability to infer patterns and peform tasks based solely on prompts, without modifying its internal parameters. The user defines the task and provides examples within the prompt. The model uses attention mechanisms to analyze these examples and deduce the task structure. The model then applies the inferred pattern to generate outputs for new inputs. For example, LargeML can use in-context learning for network intrusion detection, adaptive response to threats, or for resource management, optimizing resource allocation in real time.
Chain of thought: This approach improves reasoning by breaking complex tasks into intermediate steps. It guides the model to generate rationales for each step, processing information logically and maintaining contextual coherence. Consequently, each decision is based on established information. A practical application is fault diagnosis and troubleshooting, where LargeML analyzes network issues step-by-step, accelerating resolution and improving system resilience.
Prompt-based learning: This method uses the model’s pre-trained knowledge to generate accurate, contextually appropriate responses without additional training or parameter adjustments. Prompts focus the model’s attention and shape its output for specific tasks. For example, LargeML can analyze historical data to predict maintenance needs and preempt outages by identifying failure patterns.
II-B2 LargeML Operation
LargeML operates through a multi-stage process, beginning with comprehensive data collection and preprocessing to ensure quality. It undergoes initial training on advanced algorithms and high-performance hardware to identify data patterns, followed by fine-tuning for specific tasks to enhance accuracy. Post-deployment, LargeML processes real-time user inputs, delivering precise responses while continuously adapting to feedback and new data, maintaining its accuracy and relevance over time.
However, LargeML operation varies with the application domain and character of input data from sensors and devices [5]. In wireless communication, LargeML is typically used to optimize connectivity and manage network traffic across mobile networks, Wi-Fi systems, and satellite links. In IoT contexts, LargeML analyzes data from smart home devices, connected vehicles, and healthcare monitors to support intelligent automation and real-time decisions. In smart homes, LargeML functions as an AI control platform, adjusting lighting, temperature, and security systems based on user behaviors. In connected vehicles, LargeML enables real-time navigation, vehicle diagnostics, and autonomous driving by processing sensors and wireless network data. In healthcare, LargeML supports remote monitoring of vital signs, enabling prompt interventions and personalized care.
LargeML can also be deployed on edge devices and cloud platforms for efficient system management [52]. Edge deployment reduces latency by enabling local data processing, while cloud integration offers the scalability required to manage large volumes of data and user interactions. Continuous monitoring and updates based on user feedback and new data help maintain the model’s accuracy and relevance. Such integration across wireless and IoT systems enables more responsive and efficient solutions in a range of domains. For example, LargeML can optimize traffic flow and energy consumption in smart city infrastructure, enable predictive maintenance through industrial IoT monitoring, and improve agricultural practices based on real-time environmental data.
II-B3 Performance Evaluation
Given its broad application scope, the performance of LargeML can be evaluated at three levels. At the learning level, key indicators include model accuracy, output quality, and operational efficiency. Model accuracy is measured using metrics [55] such as precision, recall, F1 score, and area under the receiver operating characteristic (AUC) curve, which assesses the model’s ability to correctly classify inputs. For example, AUC measures the model’s ability to discriminate between classes. Model output quality is assessed using metrics such as perplexity, bilingual evaluation understudy (BLEU) score, and recall-oriented understudy for gisting evaluation (ROUGE). Perplexity gauges the model’s confidence in predicting subsequent outputs. BLEU and ROUGE compare the model’s generated output to reference resources.
At the operational level, performance is measured by efficiency metrics, such as latency, throughput, and resource use. Latency refers to the time required for the model to generate a response. Throughput indicates the number of requests the model can process per unit of time. Resource use tracks the computational resources consumed to ensure the model operates efficiently without overloading the system.
At the service level, evaluation focuses on user experience, ethics, and accountability [17, 19]. User experience is assessed by: (i) User satisfaction to capture user perceptions of model performance via direct feedback, (ii) Engagement to measures the frequency and depth of user interactions with the model, and (iii) Retention rate to indicate the proportion of users who continue using the model over time. Meanwhile, ethics and accountability metrics assess whether LargeML meets critical standards: (i) Fairness to evaluates whether the model’s predictions are fair across different user groups, helping identify and reduce bias, (ii) Transparency to measure how clearly users and stakeholders can understand the model’s decisions, often supported by explainability tools, and (iii) Ethical compliance to ensures adherence to privacy, security, and accountability standards, fostering trust in the system.
III Motivations and Requirements for TinyML– LargeML Integration in 6G Networks
III-A Challenges in TinyML
III-A1 Resource Constraints
TinyML devices, such as edge IoT nodes equipped with sensors or microcontrollers, operate under stringent processing power, memory, and energy constraints that complicate the deployment of complex ML models, forcing trade-offs between model accuracy and device capabilities. The resource constraints for TinyML device vary. For example, as noted in [56], microcontrollers may lack DRAM and an operating system and operate with up to 1 MB of read-only flash memory and up to 256 KB of SRAM. The study in [11] compares hardware platforms that support TinyML devices, reporting flash memory capacities of up to to 128 MB and SRAM capacities up to 64 MB. Consequently, given the absence of sufficient CPU, GPU, or TPU resources, the deployment and training of ML models on such devices is impractical. Larger models, although more accurate, often exceed the memory capacities of TinyML devices, and require aggressive model compression or simplification. In the context of 6G, such limitations become more pronounced as the demand for real-time, data-intensive applications increases.
III-A2 Model Compression
The deployment of ML models on TinyML devices requires compression techniques such as pruning, quantization, and knowledge distillation to reduce model size and complexity [57]. However, these methods often degrade model accuracy and performance, introducing the need for a careful trade-off between compression efficiency and model integrity. For example, quantized graph optimization presents particular challenges due to mixed-precision tensors and the absence of batch normalization layers [58]. In addition, TinyML devices lack the computational capacity to support full backpropagation. Although pruning reduces model size by eliminating less critical parameters, it may also inadvertently remove essential parameters, resulting in lower prediction accuracy.
III-A3 Energy Efficiency
Many TinyML devices operate on battery power, requiring models with high energy efficiency to ensure sustained operation. The design of TinyML models that minimize power consumption while satisfying performance requirements is a challenging task, given the computational overhead of ML inference. Poor energy efficiency can quickly deplete power sources, shorten device lifespan, and limit deployment in remote or inaccessible locations. For example, continuous data processing on a battery-powered IoT sensor can rapidly exhaust the power supply if the model lacks an energy optimization mechanism [40]. In real-time applications, such as autonomous vehicles or industrial automation, which are key areas in the 6G paradigm, low-latency inference is critical. Achieving fast inference without sacrificing energy efficiency requires advanced optimization strategies.
III-A4 Data Preprocessing and Limited Memory Resources
TinyML devices face significant constraints in performing on-device data preprocessing due to limited computational and memory resources. Tasks such as normalization, augmentation, and feature extraction require substantial processing power and memory, which TinyML devices generally lack. For example, image data captured by a camera may need extensive preprocessing before it can be used for object recognition [59]. Reduction of the dataset to satisfy memory constraints can also impair model robustness and reliability.
III-A5 Mission-Critical Reliability
TinyML’s limited numerical precision and potential for error during circuit fabrication and wafer-level assembly raise concerns about system reliability. These concerns are amplified in mission-critical applications. For example, in healthcare, TinyML may be used for remote surgery, vital sign monitoring, diagnosis, or treatment management, where any error could lead to severe or life-threatening consequences [7]. Reliability and robustness are therefore non-negotiable in such contexts, where device failures could directly affect patient safety.
III-A6 Data Privacy and Security
TinyML frequently processes data from sensors such as cameras and microphones, which may inadvertently capture sensitive personal information. This raises serious privacy concerns, especially as 6G is expected to interconnect billions of IoT devices[11]. Implementation of privacy-preserving techniques, such as differential privacy or secure multiparty computation, is challenging on TinyML platforms due to their limited computational capacity. Such limitations also increase vulnerability to unauthorized manipulation or access by malicious actors [60].
III-B Challenges in LargeML
III-B1 High Resource Consumption
LargeML models, which contain millions to trillions of parameters [61], require substantial computational resources for both training and inference. These models typically depend on high performance GPUs or specialized hardware such as TPUs. Training LargeML can span days or weeks, consuming considerable energy and raising environmental concerns. Their size, along with huge datasets, necessitates extensive memory and storage capacity. High-capacity RAM and storage systems are essential equipment for managing weights and activation states during both training and inference.
III-B2 Latency Sensitivity and Real-Time Processing
The complexity and size of LargeML models result in increased inference latency. In addition to execution time, these models are often deployed on cloud servers to meet their resource demands, requiring frequent data transfers between edge devices and the cloud. This dependence introduces network latency, which can be inconsistent and difficult to control. Many 6G applications demand real-time or near-real-time responsiveness, requirements that LargeML may struggle to meet. For example, 6G real-time applications such as augmented reality, virtual reality, and autonomous driving are highly sensitive to latency and cannot accommodate the delay caused by both inference processing and data transmission. [62].
III-B3 Deployment and Accessibility
The deployment of LargeML involves a series of complex tasks, including hardware setup, software configuration, and model optimization for specific applications. These steps require substantial technical expertise and financial investment. Hardware, such as GPUs, TPUs, and high-capacity servers, is often prohibitively expensive. For advanced LargeML models designed to achieve high accuracy on complex tasks, the costs can further restrict accessibility, reproducibility, and broader adoption, particularly among small and medium-sized enterprises and independent researchers [63]. In many developing regions, infrastructure to support LargeML remains inadequate, hindered by unreliable internet access, limited data centers, and a shortage of skilled professionals. These disparities widen the digital divide, limiting the reach of AI advances as the 6G era approaches.
III-B4 Scalability
The scaling of LargeML presents significant challenges due to high computational costs, data management complexities, and stringent network requirements [64]. Resource demands frequently exceed the capabilities of current infrastructure. LargeML requires extensive data for training and continuous updates, with consequent resource-intensive, complex data management, compounded by heterogeneous datasets sourced from a range of domains served by 6G networks. Managing such datasets is resource-intensive, and inefficient handling leads to data bottlenecks and degraded model performance. Integration of heterogeneous data sources adds further complexity, requiring extensive preprocessing and validation. As model scale increases, supporting infrastructure must enable automation and orchestration to manage higher data throughput and maintain low-latency communication between distributed systems. Network constraints can become critical bottlenecks, undermining scalability and compromising the real-time performance demanded by 6G applications.
III-B5 Interpretability and Fine-Tuning
LargeML models often function as “closed boxes”, with internal decision-making processes that lack transparency. The complexity of interactions between millions of parameters and layers obscures how specific decisions or predictions are generated. This opacity is unacceptable in mission-critical 6G applications such as healthcare, autonomous vehicles, and finance, where trust and accountability require explainable outputs. Regulatory requirements, for example the European Union’s General Data Protection Regulation [65], explicitly mandate the right to explanation, further reinforcing the need for interpretable AI. Effective model adaptation also depends on fine-tuning that is representative of the target data domain [66]. Fine-tuning involves adjustment of learning rates and batch sizes, regularization of parameters, or retraining of model components. This process must balance performance gains against the risks of overfitting or underfitting problems. Inconsistencies between training and data fine-tuning degrade model accuracy and limit real-world applicability.
III-B6 Data Privacy and Security
LargeML models typically aggregate data from multiple sources into centralized databases to support server-based training and inference. However, such centralization introduces significant privacy risks, as large repositories of sensitive information become prime targets for cyberattack [67]. Even during inference, LargeML models may requires access to confidential data, complicating privacy protection in real time, particularly when deployed across distributed networks. In addition to data privacy, securing LargeML against threats such as model extraction [68] and adversarial attacks [69] remains critical. Attacks can compromise model integrity and expose sensitive training data. For example, an adversary may reverse-engineer a model to recover private information embedded in its training data.
III-C The Vision for 6G: Objective Requirements and the Assistance of Bidirectional Integration
Figure 3: Envisaged 6G capabilities and corresponding pivotal general requirements.
IMT-2030 (6G) will support 15 key capabilities: nine enhanced from 5G (e.g., reliability, latency, data rates) and six new ones (e.g., coverage, sensing, AI), as per [70]. Some metrics are qualitatively described due to complexity [71]. These align with six 6G requirements: ubiquitous connectivity, extreme performance, high intelligence, strong security, green communication, and efficient ISAC (see Fig. 3).
Standalone technologies cannot meet 6G demands (TinyML and LargeML limitations are evident in Sections III-A, III-B), and thus, the focus will shift to an open wireless architecture, with integrations of multiple key technologies from 4G and beyond-5G [72]. Bidirectional TinyML–LargeML integration leverages 6G infrastructure: TinyML processes data at the edge, sending critical information to LargeML for advanced analytics and decision-making. Specific approaches (TL, FTL, SL, FSL) are detailed in Section IV.
III-C1 Ubiquitous Connectivity
6G will accommodate the exponential growth of interconnected devices, including IoT, sensors, wearables, and smart infrastructure, via scalable architecture and dynamic resource management. It must support connection densities of $10^{6}-10^{8}$ devices/km2 and data volumes five times greater than 5G [73]. Key enablers include network slicing and virtualization [74], which allow efficient resource allocation across diverse devices. Deep integration with AI and ML frameworks will enable pervasive, intelligent functionality. A promising approach involves the bidirectional integration of TinyML and LargeML. TinyML handles local data processing at the edge, reducing latency and offloading servers, while LargeML, deployed across a ubiquitous server set (cloud, edge, mobile devices) [75], provides a global view for system-wide optimization. This approach supports decentralized computing and further enhances scalability.
Beyond scalability, 6G aims to deliver ubiquitous global coverage, including terrestrial, aerial, satellite, and underwater systems [76, 77], ensuring connectivity in remote, challenging environments. It must also support high-mobility platforms ($500-1000$ km/h), requiring seamless handovers and latency management strategies [78]. AI-powered techniques like mobility prediction and intelligent handover strategies are expected to ensure seamless and reliable connectivity [79]. The integration of TinyML and LargeML is central to this approach. TinyML enables real-time monitoring and handover prediction at the device level, while LargeML aggregates data across the network to refine mobility and handover strategies. Despite the heterogeneity and non-independent and identically distributed (non-IID) nature of data generated by edge devices [80], LargeML frameworks can facilitate collaborative learning and efficient data aggregation. For example, wearable devices equipped with TinyML can monitor patient vitals and detect anomalies locally. Given the non-IID nature of health data, from among patients and over time from a single patient, LargeML performs server-side aggregation and analysis to support personalized healthcare delivery.
III-C2 Extreme Performance Networking
6G networks require extreme performance across multiple dimensions, including ultra-high peak and user-experienced data rates, ultra-low latency, and high reliability. Quantitatively, reliability targets range from $99.999$ to $99.99999$ with latency goals between $0.1$ and $1$ ms. As outlined in [81], other metrics are qualitative. Data rate expectations for 6G far exceed those of 5G. Peak data rates are projected between $50$ to $200$ Gbps, potentially reaching $1$ Tbps, while user-experienced rates are expected to range from $300$ to $500$ Mbps.
Integrating TinyML and LargeML enhances 6G systems by combining real-time, lightweight local processing with advanced server-side computation. TinyML reduces latency and bandwidth use for simple tasks, while LargeML handles complex computations. In the 6G environment, where servers are ubiquitous, exchanging trained models instead of raw data lowers communication latency, and hierarchical model distribution improves response times and throughput. Bidirectional integration enables collaborative learning through model splitting and transfer, optimizing distributed computing for both edge and server training. From a reliability perspective, integrating TinyML and LargeML harnesses the powerful capabilities of LargeML for server-based decision-making in mission-critical applications. These models can be fine-tuned in real time using data from TinyML devices, ensuring consistent performance and reliability. Integration allows multiple devices to jointly validate shared data without compromising privacy or security. Redundancy in this manner strengthens model accuracy and resilience to data loss or corruption.
III-C3 High Intelligence
The 5G era has evolved alongside rapid developments in AI, propelling progress toward autonomous machine intelligence. With the advent of 6G, a new phase of distributed intelligence is emerging, marked by seamless integration of distributed data processing, learning, computing, and AI inference across large-scale, interconnected systems [81]. This network of smart devices is strengthened by the complementary roles of TinyML and LargeML. TinyML enables edge devices to perform local data processing and training, transmitting only essential outputs to servers, thereby enabling real-time decisions. LargeML, operating on powerful servers, handles resource-intensive tasks, optimizing service intelligence and personalization through a decentralized architecture. Besides, the bandwidth demands of 6G use cases are substantial, for instance, high-accuracy ISAC requires 0.75 GHz, while holographic communications may need up to 1.1 GHz [73]. To meet these needs, AI-driven spectrum optimization will be vital. Real-time AI monitoring can guide dynamic spectrum allocation, reducing congestion and enhancing frequency utilization. AI-empowered cognitive radio networks will enable devices to sense and adapt to their environments, while AI-enabled signal processing will improve signal quality and spectrum efficiency. AI will also facilitate rapid interference mitigation, spectrum sharing, and reallocation of underused bands, ensuring optimal network performance. The convergence of AI and 6G is poised to drive innovation across domains such as smart cities, autonomous transport, healthcare, and immersive entertainment. AI-integrated 6G networks will support real-time urban infrastructure management, enhance the safety and responsiveness of autonomous vehicles, expand access to AI-powered medical diagnostics and remote care, and deliver deeply interactive, high-fidelity entertainment experiences.
III-C4 High Privacy and Security
Privacy and security are critical in deploying ubiquitous AI tasks over 6G networks [82]. As the volume of sensitive data grows, protecting it from unauthorized access and addressing ethical concerns, such as transparency in AI decision-making and privacy risks, becomes imperative. Increasing interconnectivity and system complexity heighten vulnerability to cyber threats, necessitating robust security frameworks that comply with evolving data protection regulations without compromising performance.
While the integration of TinyML and LargeML helps mitigate some privacy concerns, further technologies are needed to address security comprehensively. TinyML enables local preprocessing at the edge, extracting essential features and minimizing the transmission of sensitive data, thus preserving user privacy. It also supports collaborative model training without sharing raw data, reducing breach risks. However, TinyML’s limited resources may restrict implementation of advanced privacy-preserving techniques. Centralized and distributed data repositories remain key targets for attacks in a hyperconnected 6G landscape. TinyML further enhances resilience by enabling real-time anomaly detection at the edge, allowing for rapid threat identification and response. In tandem, LargeML systems leverage high-performance servers to perform in-depth threat analysis using aggregated data from multiple devices. These systems can deploy adaptive, sophisticated security mechanisms, as discussed in [83], bolstering the overall security posture of 6G networks.
III-C5 Green Communications
Green communications prioritize energy efficiency and sustainability, core elements of the 6G vision aimed at reducing CO2 emissions and minimizing the environmental impact of communication technologies [84]. As 6G expands the number of connected devices, energy consumption is expected to rise, making the reduction of the information and communication technology sector’s carbon footprint a critical goal. Energy-efficient systems not only cut operational costs but also benefit both providers and users. The combined use of TinyML and LargeML advances this goal by optimizing resource allocation and energy efficiency across the network. TinyML enables low-power devices to process data locally, reducing transmission frequency and reliance on energy-intensive servers. LargeML complements this by efficiently managing data at the server level, leveraging its computational capacity without unnecessary overhead. Their integration distributes computation across edge and cloud environments, alleviating the load on central servers and data centers, thereby significant reducing CO2 emissions.
TinyML further supports sustainable operations by adapting to real-time energy constraints and incorporating energy harvesting (e.g., solar or kinetic sources), minimizing dependence on traditional power. When paired with LargeML, these systems maintain performance even under tight energy budgets. In environmental monitoring, TinyML enables on-site anomaly detection, forwarding only critical data to centralized systems for deeper analysis, facilitating timely, informed responses. Additionally, their bidirectional integration supports predictive maintenance of network infrastructure, identifying issues early to prevent energy inefficiencies and environmental degradation. Such proactive measures are vital to reducing the ecological footprint of future communication networks.
III-C6 Efficient ISAC
ISAC is a key 6G requirement, unifying advanced sensing, communication, and precise positioning within a unified framework [85]. ISAC supports critical functions such as range, velocity, and angle estimation, as well as object detection, localization, imaging, and mapping. This integration reduces redundancy, optimizes resource use, and boosts overall network performance. High-precision sensing is essential for latency-sensitive applications like autonomous vehicles, smart cities, and AR, all of which demand real-time, location-specific data. Positioning accuracy in 6G is targeted within $1-10$ cm [81]. The integration of TinyML and LargeML can significantly enhance the efficiency, accuracy, and robustness of ISAC systems. For example, TinyML enables on-device processing and feature extraction, allowing initial estimation of key parameters and basic object detection. This local filtering reduces noise and improves signal fidelity. Processed outputs are then sent to LargeML systems for deeper analysis and validation, improving detection accuracy and reducing false positives. Additionally, TinyML supports real-time mapping by continuously updating environmental models, which is crucial for autonomous navigation that relies on high-resolution, up-to-date spatial information.
IV Efficient bidirectional Integration Solutions for 6G and Beyond Networks
IV-A Transfer Learning
Figure 4: TinyML–LargeML system with TL.
TL, a subset of ML methods, focuses on reusing the knowledge from pre-trained models across related tasks and data domains. The technique provides an effective approach to TinyML–LargeML integration in 6G and beyond networks by combining the strengths of server-based large-scale training and on-device adaptability [86]. Fig. 4 depicts a TL-based two-way integration framework, consisting of three main steps.
1
Server-side training: A large dataset of $D_{S}$ samples, denoted ${{\cal D}_{S}}=\left\{{1,\ldots,{D_{S}}}\right\}$, is used to train a LargeML model, denoted $f_{S}$, on a high-performance cloud or edge server, referred to as the source domain. Let ${\bf{\Theta}}=\left\{{{{\bf{w}}_{l}},{{\bf{b}}_{l}}}\right\}_{l=1}^{L}$ be the set of parameters comprising all weights ${{{\bf{w}}_{l}}}$ and biases ${{{\bf{b}}_{l}}}$ across $L$ layers, which form the model pipeline. The training objective is to minimize the loss function, i.e.,
${\bf{\Theta}}=\arg\mathop{\min}\limits_{\bf{\Theta}}\frac{1}{{{D_{S}}}}\sum\nolimits_{i\in{{\cal D}_{S}}}{{\cal L}\left({{{\bf{y}}_{i}},{f_{S}}\left({{{\bf{x}}_{i}}}\right)}\right)},$ (1)
where ${\left\{{{{\bf{x}}_{i}},{{\bf{y}}_{i}}}\right\}_{i\in{{\cal D}_{S}}}}$ represents the training samples, and ${{f_{S}}\left({{{\bf{x}}_{i}}}\right)}$ denotes the model’s activation after processing input feature ${\bf{x}}_{i}$. The loss function ${\cal L}(\cdot)$ varies depending on the learning task and may include metrics such as cross-entropy, mean squared error, MAE, or log-likelihood. For example, the loss function of a linear regression problem over dataset $\left({{\bf{X}},{\bf{y}}}\right)$ is given by ${{\cal L}_{0}}\left({{\bf{w}}|{\bf{X}},{\bf{y}}}\right)=\frac{1}{M}\sum\nolimits_{m=1}^{M}{{{\left({{{\bf{w}}^{T}}{{\bf{x}}_{m}}-{y_{m}}}\right)}^{2}}}$, where ${{{\bf{x}}_{m}}}$ and $y_{m}$ are the input feature vector and label scalar, respectively, of the data sample $m$ among $M$ data points of $\left({{\bf{X}},{\bf{y}}}\right)$. The server’s high computational resources and storage capabilities allow it to process large datasets and train complex models effectively.
Server-side training can adopt either a standalone architecture (e.g., two-tiered client-edge, two-tiered client-cloud) or a multi-level collaboration model [87]. Compared to standalone setups, multi-level collaboration offers greater flexibility in balancing communication and computational overheads. Four representative scenarios include: (i) Hybrid training, where initial training of LargeML is conducted in the cloud, followed by deployment to edge servers for fine-tuning and real-time adaptation using local data; (ii) Distributed fashion, where the training workload is distributed between the cloud and edge servers, with large-scale, resource-intensive tasks handled in the cloud and latency-sensitive, less intensive tasks processed at the edge; (iii) Hierarchical processing, where edge servers perform data preprocessing to reduce data volume and complexity before transmitting it to the cloud for deeper analysis, while updates from the cloud can subsequently relayed back to edge servers for local execution and adaptation; and (iv) Enhanced availability, wherein hierarchical server models improve system resilience by allowing edge servers to operate independently during cloud outages.
2
Pre-trained knowledge transfer: After training on the source domain, selected components of the LargeML $f_{S}$, specifically a subset of ${\bf{\Theta}}$, are transferred to the target domain (i.e., TinyML devices). These transferred components encapsulate the general features and patterns learned from ${\mathcal{D}}_{S}$. Effective knowledge transfer in TL hinges on four primary considerations as described below [88, 89, 90].
What to transfer
The “what” aspect refers to identifying the source model elements to be reused in the target domain. These may include the components or knowledge that are likely to benefit the target task. We denote $t_{S}$ as the transferred model, where $t_{S}$ is a subset of ${\bf{\Theta}}$. This decision is the most critical in TL.
When to transfer
The “when” aspect involves determining the appropriate timing and frequency of knowledge transfer between the source and target domains. It requires the selection of a strategy for updating the target domain with knowledge from the source domain. Strategies can be grouped into three types [91]: (i) Real-time TL: Knowledge transfers occur immediately after updates in the source domain, which is suitable for dynamic environments requiring rapid adaptation; (ii) Non-real-time or semi-real-time TL: Knowledge transfers follow a fixed schedule (e.g., daily, weekly) or occur at defined operational checkpoints, which suits stable environments or models where immediate updates are not critical; and (iii) On-demand TL: Knowledge transfers occur only when explicitly requested by the target domain, which benefits scenarios in which the target domain operates autonomously and requires updates under specific conditions only.
Where and how to transfer
The “where” aspect refers to identifying the target domain where the knowledge will be applied (e.g., TinyML devices). The “how” aspect ensures compatibility and maximizes the effectiveness of knowledge use in the target domain. Effective transfer depends on both domain similarity and the selection of optimization techniques. Transfer is typically more successful between domains with overlapping characteristics. For example, a resource management model trained on 5G network infrastructure may be adapted to monitor and optimize components in a 6G network [92]. The hardware and software constraints of TinyML devices also affect both where and how knowledge transfer occurs. The transferred model must be compatible with these constraints to function efficiently under limited computational and memory resources.
Examined through these guiding questions, TL can be categorized according to several criteria [89]. One approach focuses on label availability: (i) In transductive TL, labeled data is available only in the source domain, and the source and target tasks are similar; (ii) In inductive TL, labeled data is available only in the target domain, and the source and target tasks differ; and (iii) In Unsupervised TL, neither domain has labeled data, and the target task is distinct yet related to the source task. Another categorization approach depends on the alignment of feature and label spaces between the source and target domains. When both domains share identical features and label spaces, the setting is classified as homogeneous TL; otherwise, it is considered non-homogeneous TL. A third, widely used classification is based on the transfer mechanism. Specifically, Instance-based TL selects and re-weights specific instances (data points) from the source domain to improve their relevance in the target domain, aiming to identify the most useful source data points for training in the target domain [93]. Feature-based TL seeks to identify shared feature representations for the source and target domains [94]. By transforming features from both domains into a shared feature space, using techniques such as dimensionality reduction, feature extraction, or transformations, this approach facilitates more effective knowledge transfer. Parameter-based TL reuses model parameters (e.g., weights and biases) learned from the source domain and transfers them to the target domain [95]. Typically, a model is trained on a large source-domain dataset, and the learned parameters serve as an initialization point for training on the target dataset, enabling the model to retain useful knowledge while adapting to the new domain. Finally, Relational-based TL is used when data is structured as relationships or graphs (i.e., learned rules) [96]. It transfers structural or relational information, such as node embeddings, relational patterns, or graph kernels, from the source domain to support learning in the target domain.
3
Model personalization and on-device adaptation: Consider a set of $N$ TinyML devices in the target domain, denoted ${\cal N}=\left\{{1,\ldots,N}\right\}$. Each device ${n\in{\cal N}}$, equipped with limited computational resources and storage, adapts a pre-trained model $t_{S}$ along with its TinyML model $g_{n}$, using its individual, smaller, scenario-specific dataset ${{\cal D}_{n}}=\left\{{1,\ldots,{D_{n}}}\right\}$. This adaptation involves fine-tuning (i.e., freezing $t_{S}$ while updating the parameters associated with $g_{n}$) and additional training to tailor the model for relevance to the device’s specific tasks. This process is referred to as model personalization. The concept of an adaptation layer introduced in [97] augments this process by minimizing the general loss function and the maximum mean discrepancy (MMD) loss. MMD measures the divergence between source and target domain distributions by evaluating their distance in a reproducing kernel Hilbert space (RKHS). It serves as a criterion for comparing distributional discrepancy in transfer learning. An empirical estimate of the MMD loss function between ${\cal D}_{S}$ and ${\cal D}_{n}$ is given by ${{\cal M}_{{{\cal D}_{S}},{{\cal D}_{n}}}}=\big{\|}\frac{1}{{{D_{S}}}}\sum\nolimits_{i\in{{\cal D}_{S}}}{\phi\left({{{\bf{x}}_{i}},{{\bf{y}}_{i}}}\right)}-\frac{1}{{{D_{n}}}}\sum\nolimits_{j\in{{\cal D}_{n}}}{\phi\left({{\bf{x}}_{j}^{n},{\bf{y}}_{j}^{n}}\right)}\big{\|}_{\cal H}$ [97], where $\phi\left(\cdot\right)$ is the nonlinear mapping into RKHS $\cal H$, and ${\left\{{{\bf{x}}_{j}^{n},{\bf{y}}_{j}^{n}}\right\}_{j\in{{\cal D}_{n}}}}$ represents the samples from ${\cal D}_{n}$. Consequently, the objective of minimizing the loss function for local training on client ${\left\{n\right\}_{n\in{\cal N}}}$ is expressed as
${{\bf{\Theta}}_{n}}=\arg\mathop{\min}\limits_{{{\bf{\Theta}}_{n}}}\frac{1}{{{D_{n}}}}\sum\limits_{j\in{{\cal D}_{n}}}{{\cal L}\left({{\bf{y}}_{j}^{n},{t_{S}}\left({{\bf{x}}_{j}^{n}}\right)}\right)}+\lambda{\cal M}_{{{\cal D}_{S}},{{\cal D}_{n}}}^{2},$ (2)
where $\lambda>0$ is the trade-off parameter. However, given the simplicity of this shallow model, its performance is limited. To further mitigate domain shift between the source and target domains, correlation alignment (CORAL) can be applied [98]. The CORAL loss between ${\cal D}_{S}$ and ${\cal D}_{n}$ is formally defined as ${\ell_{{{\cal D}_{S}},{{\cal D}_{n}}}}=\frac{1}{{4{d^{2}}}}\left\|{{{\bf{R}}_{{{\cal D}_{S}}}}-{{\bf{R}}_{{{\cal D}_{n}}}}}\right\|_{F}^{2}$, where $d$ is the dimension of embedding features, ${\left\|\cdot\right\|_{F}}$ denotes the Frobenius norm, and ${{\bf{R}}_{\cal D}}$ denotes the covariance matrix of dataset $\cal D$. Consequently, (2) can be reformulated as
${{\bf{\Theta}}_{n}}=\arg\mathop{\min}\limits_{{{\bf{\Theta}}_{n}}}\frac{1}{{{D_{n}}}}\sum\limits_{j\in{{\cal D}_{n}}}{{\cal L}\left({{\bf{y}}_{j}^{n},{t_{S}}\left({{\bf{x}}_{j}^{n}}\right)}\right)}+\lambda{\ell_{{{\cal D}_{S}},{{\cal D}_{n}}}}.$ (3)
Since TinyML devices can perform tasks locally after knowledge transfer,they retain sensitive data on-device, eliminating the need for continuous cloud transmission or inter-device sharing, thus improving privacy.
Recent studies have explored the integration of TinyML and LargeML through TL for various applications [38, 36, 99, 100, 101, 102, 103]. For example, [38] proposes a TL-based system which combines TinyML and LargeML, named TinyTL. The framework employs ProxylessNAS-Mobile as the backbone NN for source-domain pre-training, and fine-tunes the model on TinyML devices using a pre-trained once-for-all network. The parameter-based TL approach transfers feature extractor weights. Target tasks include object and facial attribute classifications, such as recognition of cars, aircraft, flowers, birds, pets, food, and celebrities. [36] introduces a bidirectional integration framework incorporating parameter-based TL. The system pre-trains a CNN in the source domain, then adapts the CNN to TinyML devices using Tensorflow Lite Micro with pruning and quantization techniques. Target tasks include presence detection and classification inside cars. In [99], the authors developed a bidirectional integration system using feature-based TL aimed at object classification and human activity recognition for IoT health tracker applications. The source domain backbone is a deep neural network (DNN), and the target model runs on CMSIS-NN.
Building on prior work, the study in [100] proposes a feature-based TL system that exploits VGG-16, DenseNet, and MobileNet as backbone architectures for source-domain pre-training. For deployment in the TinyML domain, the system uses the EON Tuner tool, integrated with the Edge Impulse platform, to perform face mask detection and compare performance across the three models. Another study [101] applies a parameter-based TL approach for soil humidity prediction in smart agriculture. The model combines a DNN and long short-term memory (LSTM) architecture for source-domain pre-training. In the target domain, fine-tuning with TensorFlow Lite Micro and quantization techniques are used for deployment on TinyML devices. In [102], the authors introduce a feature-based TL bidirectional integration system for motion classification. The approach uses principal component analysis for dimensionality reduction and trains a CNN-LSTM model on large-scale, diverse source-domain datasets. Fine-tuning is performed on TinyML devices, with a Raspberry Pi as an edge intermediary platform. A feature-based TL system presented in [103] incorporates meta-learning for a TinyML–LargeML framework and cross-domain classification tasks. The system uses a DNN pre-trained with MCUNet, MobileNetV2, and ProxylessNAS architectures in the source domain and performs fine-tuning on TinyML devices using a task-adaptive sparse-update method. The design prioritizes adaptability to dynamic user contexts and device constraints, using few-shot learning with minimal samples.
IV-B Federated Transfer Learning
Figure 5: TinyML–LargeML system with FTL.
IV-B1 Vanilla Federated Transfer Learning
The TL strategies described in Section IV-A are effective when source and target domains share related tasks and datasets. However, their applicability in 6G environments is limited due to the prevalence of highly non-IID data across massively connected devices. FTL, which combines the strengths of FL and TL [104], is better suited for such heterogeneous settings. In 6G’s device-dense landscape, effective personalization is crucial for optimizing performance on device-specific tasks. Notably, FL enables collaborative training across distributed TinyML systems while preserving data locality on resource-constrained devices [105]. FTL enhances this by distinguishing between source clients (who participate in training the initial global model) and target clients (who personalize the model to their specific needs). As both FL and TL strategies emphasize privacy, FTL effectively addresses the dual challenges of data heterogeneity and privacy in integrated TinyML–LargeML systems. Fig. 5 illustrates the workflow of this FTL-based approach, as outlined below.
The FTL process begins with the standard FL protocol, differing primarily in the initialization of the global model. Specifically, the FL phase includes the following steps.
  • 1
    Initial global model sharing: In contrast to conventional FL, where the server initializes the global model with random weights, FTL initializes the global model using a pre-trained LargeML model $f_{S}$, trained on the server’s large dataset ${\cal D}_{S}$. As with (1), the objective function for training is given by $f_{S}^{\left(0\right)}=\arg\mathop{\min}\nolimits_{{f_{S}}}\frac{1}{{{D_{S}}}}\sum\nolimits_{i\in{{\cal D}_{S}}}{{\cal L}\left({{{\bf{y}}_{i}},{f_{S}}\left({{{\bf{x}}_{i}}}\right)}\right)}$. Once trained, $f_{S}^{\left(0\right)}$ is distributed to a source client set ${\cal K}=\left\{{1,\ldots,K}\right\}$.
  • 2
    Local training: In the training round $t$, each source client ${k\in{\cal K}}$ uses its local dataset ${{\cal D}_{k}}=\left\{{1,\ldots,{D_{k}}}\right\}$ to train and update a local copy of the model. Local training enables the model to learn patterns and insights specific to each client’s data. Mathematically, the objective for minimizing the loss function during client $k$’s local training is expressed as $f_{S,k}^{\left(t\right)}=\arg\mathop{\min}\nolimits_{f_{S}^{\left({t-1}\right)}}\frac{1}{{{D_{k}}}}\sum\nolimits_{j\in{{\cal D}_{k}}}{{\cal L}\big{(}{{\bf{y}}_{j}^{k},f_{S}^{\left({t-1}\right)}\left({{\bf{x}}_{j}^{k}}\right)}\big{)}}$.
  • 3
    Local model update: After local training, each source client uploads its updated model parameters $f_{S}^{\left(t\right)}$ to the server for aggregation.
  • 4
    Server-side aggregation: The server obtains global model parameters by minimizing the global loss function $f_{S}^{\left(t\right)}=\arg\mathop{\min}\nolimits_{f_{S,k}^{\left(t\right)}}\sum\nolimits_{k\in{\cal K}}{\frac{{{r_{k}}}}{{{D_{k}}}}\sum\nolimits_{j\in{{\cal D}_{k}}}{{\cal L}\big{(}{{\bf{y}}_{j}^{k},f_{S,k}^{\left(t\right)}\left({{\bf{x}}_{j}^{k}}\right)}\big{)}}}$ where ${r_{k}}={D_{k}}/\sum\nolimits_{k=1}^{K}{{D_{k}}}$ represents the ratio of client $k$’s data size to the total data size. This aggregation problem can be solved using the FedAvg algorithm [106], a proportion weighting strategy [107], or more advanced methods that account for data heterogeneity and varying client contributions.
  • 5
    Global model update: The updated global model is then sent back to the source clients.
Steps  2 to 5 are repeated until the global model reaches satisfactory performance.
Step 6 presents a knowledge transfer phase. Once the global model has been sufficiently trained and demonstrates strong generalization, the server transfers the useful components of the pre-trained global model $f_{S}$, denoted $t_{S}$, to the target TinyML devices (target clients). This transferred model, derived form the pre-trained global model, encapsulates the knowledge learned from source clients and serves as the foundation for the next phase.
Step 7 stands for model personalization and on-device adaptation. Target clients typically hold smaller datasets or domain-specific datasets that differ from those of source clients. To tailor the transferred model $t_{S}$ to local tasks, each target client performs personalization through TL. We consider the set of target TinyML devices ${\cal N}=\left\{{1,\ldots,N}\right\}$. Each target client fine-tunes $t_{S}$ with its local dataset ${\cal D}_{n}$. Depending on the similarity of the global model’s training data and the target client’s data, fine-tuning may involve updating the entire model $t_{S}$ or only a subset of layers. To support deployment in resource-constrained environments, each target client employs a TinyML model $g_{n}$. The personalized model is then trained using either objective functions (2) or (3). The adaptation process harnesses the advantages of TL, enabling target clients to attain high performance without the need to train a model from scratch. Consequently, this approach conserves computational resources and reduces training time.
Numerous recent technical publications have explored the use of FTL to enable seamless integration between TinyML and LargeML across diverse application domains [108, 109, 110, 111]. For example, the work in [108] introduces TinyFedTL, an open-sourced, parameter-based FTL framework designed for TinyML–LargeML systems. The framework pre-trains MobileNetV2 on a source domain and applies Perf-MobileNet, part of the TinyML Perf Benchmark, for personalized model adaptation on IoT edge devices. Experimental results validate the effectiveness of TinyFedTL in enhancing the quality of experience (QoE) under resource constraints, without compromising data privacy. Proposed in [109], FedHealth is a parameter-based FTL framework for wearable healthcare devices. FedHealth employs a CNN backbone network for source-domain pre-training, followed by fine-tuning during the TL phase to adapt the model to resource-constrained IoT devices. Evaluations of wearable activity recognition and auxiliary Parkinson’s disease diagnosis tasks demonstrate that FedHealth achieves high accuracy, without compromising privacy and security. A parameter-based FTL framework incorporating auxiliary classifier generative adversarial networks, termed ACGAN-FTL, is proposed in [110]. ACGAN-FTL runs personalized models on Raspberry Pi-based TinyML devices and is designed to predict the quality of pre-baked carbon anodes in industrial production. Recently, the authors in [111] demonstrated the superior performance of FTL compared to standalone FL, TL, and Tensorflow Lite-based fusion approaches in both classification and regression tasks in TinyML–LargeML systems. The TL principle applied in this framework centers on feature-based TL, using Genann [112] as the backbone neural network for source-domain pre-training, followed by fine-tuning for device-specific adaptation across multiple TinyML platforms, including Arduino WiFi Rev2, Arduino MKR1010, ESP8266, and ESP32.
From a hierarchical FTL perspective, the study in [113] introduces IoTDefender, an intrusion detection framework designed for resource-constrained IoT networks. IoTDefender employs three edge servers that run standard FL protocols on distinct IoT networks. Their local models are subsequently aggregated at a cloud server. IoTDefender uses a CNN backbone for source-domain pre-training, with subsequent fine-tuning for parameter-based TL and on-device adaptation in resource-constrained IoT devices. Expanding on the topology proposed in [113], the study in [114] proposes HFTL, a hierarchical FTL framework that incorporates additional fog servers. These fog servers aggregate models from associated edge servers before forwarding to a cloud server for further aggregation. HFTL uses VGG-16, ResNet50V2, and MobileNet as pre-trained models, followed by fine-tuning with an additional DNN for secure and efficient fault classification in additive manufacturing. The framework adopts parameter-based TL as its core approach.
IV-B2 Advanced Federated Transfer Learning
Building upon the foundations of FTL, below are efficient methods for integrating TinyML with LargeML.
Federated Meta-Learning (FML)
FML aims to achieve a higher level of generalization and adaptability by enabling models to learn new tasks rapidly through meta-knowledge acquired from diverse clients or domains [115]. Unlike standard FTL, which focuses on transferring knowledge across related tasks, FML trains models to adapt to previously unseen tasks. This “learning how to learn” capability improves robustness and broadens applicability, supporting both knowledge transfer and efficient task-specific adaptation.
TinyReptile, an FML-based TinyML–LargeML system introduced in [116], draws on concepts from FL and online learning. The system enables distributed resource-constrained IoT devices to collaboratively learn a robust RNN initialization for rapid adaptation to new target devices. The study evaluates TinyReptile on classification tasks using sine-wave, Omniglot, and keyword spotting datasets and demonstrates a simple yet efficient implementation on Raspberry Pi 4 and Cortex-M4 MCU edge hardware, using parameter-based TL. TinyMetaFed, a comparable framework, is introduced in [117]. Evaluated on similar TinyML devices, TinyMetaFed incorporates a modified algorithm that reduces energy consumption and communication overhead, accelerates convergence, and stabilizes training. Building on the capabilities of TinyReptile and TinyMetaFed, the study in [118] proposes an advanced prototype, deployed on similar TinyML testbed devices. The framework employs a CNN backbone architecture for pre-trained models and applies the MLPerf Tiny Benchmark for on-device, personalized model adaptation. The prototype demonstrated its effectiveness in three real-world TinyML applications: handwritten character image classification, keyword spotting, and presence detection in smart buildings. PMFed, proposed in [119], is an FML framework that employs a pre-trained CNN and fine-tunes the model on a Raspberry Pi 3B for personalized adaptation on TinyML devices. The framework demonstrated efficient performance in three real-world, IoT-enabled health monitoring tasks: prompt defibrillation following detection of ventricular arrhythmia, atrial fibrillation detection, and human activity recognition.
Federated Foundation Models (FFMs) and Federated Fine-Tuning (FedFT) of Foundation Models (FMs)
FFMs integrate the strengths of FMs and FL, enabling decentralized clients to collaboratively adapt pre-trained LargeML models without exchanging raw data. This approach combines the privacy advantages of FL with the TL benefits of TFMs [22]. FedFT of FMs allows each client to locally personalize a shared model while maintaining data confidentiality. To reduce computational overhead, PEFT techniques such as LoRA, adapter modules, and prompt tuning are widely applied [22].
The study in [120] introduces FedPT, a federated proxy-tuning method for LLMs on resource-constrained edge devices. By relying solely on model predictions rather than parameter access, FedPT reconstructs full-sized LoRA modules for server-side aggregation, reducing computational and communication overhead while matching FedPT performance in instruction-following tasks. However, its reliance on homogeneous LoRA hampers performance under non-IID data distributions. To address this, HetLoRA [121] uses zero-padding and rank-constrained truncation to standardize LoRA modules, enhancing convergence and efficiency in heterogeneous systems for tasks like multi-session chat dialogue and text summarization. CAFF [122] further improves accuracy and fairness in language and vision classification tasks by combining tiny transformers with a layer-wise FedFT and NN selection strategy. In [123], the authors present FedD2P, an efficient FedFT framework for FMs targeting heterogeneous, resource-constrained IoT devices. FedD2P distills aggregated knowledge into a prompt generator, allowing frozen FMs to adapt efficiently to downstream tasks such as image classification, satellite imagery analysis, and texture recognition. The work in [124] proposes FwdLLM, an efficient FedFT method for LLMs on heterogeneous, resource-constrained edge devices. Instead of applying traditional BP to compute precise gradients, FwdLLM combines BP-free training with PEFT by injecting minor self-generated perturbations into model parameters. Compared to existing PEFT benchmarks, this approach reduces memory footprint and accelerates convergence.
To further support model heterogeneity and computational efficiency, sparse architectures like MoE can be combined with PEFT. For example, FedFMSL [125] incorporates a two-stage MoE-based FedFT. The method first trains a global expert, then fine-tunes a local expert for improved personalization. To adapt to heterogeneous, resource-constrained edge environments, FedFMSL employs a sparsely-activated LoRA algorithm that progressively activates LoRA matrices based on historical performance tracked in a capability queue. This approach yields performance gains and resource savings in various image classification tasks, for example food recognition, land use, land cover, and human action detection. In [126], the authors introduce FedMoE, a sparsely-activated MoE operating with LoRA-based FedFT for heterogeneous, resource-constrained edge environments. FedMoE dynamically adjusts submodels using global expert recommendations and achieves fast convergence and improved performance in text classification, summarization, and reading comprehension tasks. Lastly, A3SMoE [48] introduces an efficient FedFT method for FMs tailored to heterogeneous, resource-restricted edge settings. A3SMoE extends the concept of data-conditional computation in sparsely-activated MoEs to resource-conditional computation, enabling clients to activate suitable experts based on resource availability. It outperforms prior LoRA-based techniques in instruction-tuning and complements existing heterogeneous LoRA methods.
IV-C Split Learning
SL is an ML technique designed for collaborative model training under resource-constrained environments, enabling full model training on devices to improve learning efficiency [75]. SL is particularly suitable for integrating TinyML and LargeML within 6G and future network architectures. By partitioning the model across client devices and servers, SL allows TinyML devices to participate in training and inference tasks without becoming overwhelmed by the computational demands of LargeML while also preserving data privacy. Two main SL variants can be applied to TinyML–LargeML systems: vanilla SL (VSL) and parallel SL (PSL).
Figure 6: TinyML–LargeML systems: (a) VSL and (b) PSL.
IV-C1 Vanilla Split Learning
VSL is the foundational form of SL in which the training process proceeds sequentially fro client to client [127]. On the server side, either a standalone server architecture (e.g., two-tiered client-edge or two-tiered client-cloud) or a multi-level server collaboration model can be implemented.
[…]
The integration of VSL-based TinyML with LargeML has enabled a diverse range of applications for tasks in resource-constrained environments. For example, a two-stage wireless channel adaptive split inference method [130] addresses memory and energy constraints on edge devices. The experimental results demonstrate enhanced privacy, improved inference accuracy, and dynamic adjustment of split points based on real-time variations in wireless channel gain. ARES [131] is an adaptive, resource-aware SL scheme designed for efficient model training in resource-constrained IoT systems. ARES mitigates the impact of stragglers by employing device-specific split layers, balances energy consumption with training time, and adapts to fluctuating network throughput and computing capacity. An adaptive SL algorithm based on Lyapunov theory is introduced in [132]. This approach dynamically selects split layers for resource-constrained devices and allocates server-side computing resources in wireless edge networks. Simulations demonstrate significant reductions in average training delay and energy consumption.
When multi-level server collaboration is employed, multiple cloud and edge servers jointly participate in VSL through a strategy known as multi-hop SL [75]. In this architecture, multiple hierarchical servers are organized in sequence, forming a serial multi-hop SL framework, as described in [133]. Here, the model is partitioned into multiple segments: (i) initial layers are deployed on TinyML devices; (ii) middle layers are distributed across multiple servers, each managing a distinct segment; (iii) final layers reside on a server that generates the output. This setup follows a similar structure to standalone server architectures but involves sequential inference and BP across multiple servers [134, 135, 133]. Specifically, the first client processes raw local data through its local model segment to generate smashed data before forwarding it to the first server. The server processes this data through its model segment and forwards the output to the next server, continuing through the chain until the final output is produced. During BP, the final server computes the loss, performs gradient updates, and propagates gradients backward through the server chain. Each intermediate server updates its own parameters and passes gradients to the preceding server or client. Clients then perform local BP and parameter updates in sequence, completing the training loop [134, 135]. In a system with $N$ TinyML devices and $K$ collaborating servers, each inner iteration requires $2NK$ inference and BP steps, highlighting the need for efficient algorithms to reduce latency.
Beyond serial data flow, centralized smashed data routing [75] can also be implemented in 6G networks by integrating ad-hoc data routing and centralized model partitioning, enabled through software-defined networking. This approach must consider bandwidth, memory, and computing constraints. Effective server coordination is crucial for compatibility between model segments, especially with non-IID data distribution across TinyML devices, necessitating periodic synchronization of model parameters for consistency.
IV-C2 Parallel Split Learning
The sequential training process of VSL introduces substantial training latency and does not scale as the number of participating TinyML devices increases, a problem that will be exacerbated by the massive connectivity expected in 6G networks. The highly non-IID nature of client data can also degrade model performance in sequential training, since the model may become biased toward the data distribution of the previously trained client. To address these challenges, PSL enables multiple TinyML devices to train their local model segments concurrently [136, 137]. This parallelization reduces total training time, introduces “semi-scalability” relative to the number of participating devices, and partially mitigates the impact of non-IID data in 6G mediums.
Fig. 6(b) illustrates the workflow of PSL-based TinyML–LargeML systems under a standalone server architecture. Unlike VSL-based bidirectional integration, which requires $N$ sequential rounds of forward and backward propagation, PSL completes training with a single round, regardless of $N$. This is the basis of semi-scalability, meaning the number of propagation rounds is scalable on the client side. However, the server-side training workload (i.e., smashed data gradients and computational workload), still scales proportionally with $N$. The inner-iteration training process of a PSL-based TinyML–LargeML system proceeds as follows:
  • 1
    Simultaneous client-side local FP: Each TinyML device ${\cal N}=\left\{{1,\ldots,N}\right\}$ trains on its local datasets ${{\cal D}_{n}}=\left\{{1,\ldots,{D_{n}}}\right\}$ using its model segment ${\bf{T}}_{n}$, generating smashed data ${\bf{T}}_{n}({\bf{x}}_{n})$ at the designated split layer.
  • 2
    FP of smashed data: Each device transmits its smashed data and corresponding labels $\left\{{{{\bf{T}}_{n}}\left({{\bf{x}}_{n}}\right),{\bf{y}}_{n}}\right\}$ to the server. The server receives these inputs concurrently from all devices.
  • 3
[…]
  • 4
    BP of smashed data gradients: The computed gradients ¯$\nabla\bar{\cal L}({\bf{T}}_{n}({\bf{x}}_{n}))$ are returned to each respective TinyML device in parallel.
  • 5
    Simultaneous client-side local BP: Each TinyML device performs local BP and updates its model parameters, i.e., ${\bf{T}}_{n}\leftarrow{\bf{T}}_{n}-\eta_{n}\nabla{\bar{\cal L}}({\bf{T}}_{n}({\bf{x}}_{n}))$.
These training steps are repeated over consecutive rounds until the model converges, i.e., minimizing the global loss function
$\mathop{\min}\limits_{{\bf{L}},{{\left\{{{{\bf{T}}_{n}}}\right\}}_{\forall n}}}\frac{{\sum\nolimits_{n\in{\cal N}}{\sum\nolimits_{j\in{{\cal D}_{n}}}{{\cal L}\left({{\bf{y}}_{j}^{n},{\bf{L}}\left({{{\bf{T}}_{n}}\left({{\bf{x}}_{j}^{n}}\right)}\right)}\right)}}}}{{\sum\nolimits_{n\in{\cal N}}{{D_{n}}}}}.$ (4)
After convergence, each TinyML device fine-tunes the model to its specific task. However, two potential issues arise:
  • Although PSL helps prevent overfitting caused by aligning the end-to-end output model too closely with the distribution of the previous client, the highly non-IID nature of data across TinyML devices can reduce the convergence rate. This may require server or client-side aggregation strategies. Relevant solutions include EdgeSplit [138] and FSL.
  • The server assumes the primary training workloads from $N$ TinyML devices, which does not scale as $N\to\infty$. Although cloud/edge servers are generally powerful, they could still become bottlenecks in the massive connectivity expected in 6G systems. An efficient variant, PSL (EPSL), addresses this issue [139].
Further discussion of EdgeSplit and EPSL is given in Section IV-C3; details of FSL are provided in Section IV-D.
Recent publications have applied PSL to TinyML–LargeML systems [140, 141, 142]. For example, the study in [140] proposes a privacy-sensitive PSL framework that supports SL training across multiple resource-constrained devices in parallel. The framework addresses overfitting caused by non-IID training orders and varying dataset sizes on resource-constrained devices while maintaining data privacy. Introduced in [141], AdaSplit represents an efficient scaling SL method for heterogeneous, resource-constrained systems. AdaSplit minimizes on-device computation and optimizes server-side processing and communication. It eliminates the need for server gradients at the client side, reduces payload size and interaction frequency to reduce communication overhead, and support updates to sparse model partitions on heterogeneous resource-constrained devices. The work in [142] formulates and solves a multiplayer bargaining problem to identify optimal split-layer strategies in PSL-based bidirectional integration systems. The approach balances computation and wireless transmission energy consumption, training time, and data privacy. The proposed framework also demonstrates robustness with non-IID datasets across resource-constrained devices.
In PSL-based TinyML–LargeML systems involving multi-level server collaboration, the architecture extends beyond simple client–edge or client–cloud configurations by incorporating multiple servers or hierarchical layers of servers that collaborate during the training process. Such setups require server orchestration and aggregation strategies, necessitating advanced PSL schemes to coordinate interactions between server layers and integrate TinyML with LargeML seamlessly. Detailed in Section IV-C3, this class of schemes is referred to as advanced PSL, where parallel-server PSL [143] and multi-hop PSL [144] serve as pioneering frameworks. These frameworks support the development of more sophisticated server orchestration and aggregation methods.
IV-C3 Advanced Parallel Split Learning
This section introduces EdgeSplit and EPSL, which are advanced PSL schemes designed to address the highly non-IID nature of data across TinyML devices. Both schemes consider model aggregation at the server side. EPSL further provides full scalability. The details of each scheme are described below.
Figure 7: Integrated TinyML–LargeML system with EdgeSplit.
EdgeSplit [138] is an advanced PSL variant designed to address the challenge of highly non-IID data distributed across TinyML devices. Non-IID data causes local models trained on individual TinyML devices to diverge, reducing generalization and degrading global model performance. EdgeSplit introduces a coordinated learning approach across diverse datasets while preserving the core benefits of PSL. Fig. 7 illustrates the inner-iteration training steps of EdgeSplit, which begins by a standard PSL procedure: 1 Simultaneous client-side local FP, 2 FP1 of smashed data, 3 server-side local FP and local BP, 4 BP1 of smashed data gradients, and 5 simultaneous client-side local BP. The subsequent steps introduce the key advancements of EdgeSplit over conventional PSL.
  • 6
    FP2 with updated weights: After client-side local BP, each TinyML device forwards its updated model weights ${\bf{T}}_{n}$ to the server. These weights incorporate the local learning adjustments made by each device based on its non-IID data.
  • 7
    Server-side aggregation: The server aggregates all updated weights received from TinyML devices using an algorithm, such as FedAvg [106]:
    ${{\bf{T}}_{{\rm{FedAvg}}}}={{\sum\nolimits_{n\in{\cal N}}{{D_{n}}{{\bf{T}}_{n}}}}}\Big{/}{{\sum\nolimits_{n\in{\cal N}}{{D_{n}}}}}.$ (5)
    This step harmonizes learning across all devices by combining their weights and enables the server to produce a more generalized model that reflects the diversity of data distributions.
  • 8
    BP2 of aggregated weights: The server broadcasts the aggregated weights ${{\bf{T}}_{{\rm{FedAvg}}}}$ to all TinyML devices, ensuring that each device updates its model with globally informed weights that reduce the impact of non-IID data.
  • 9
    Updated client-side model: Each TinyML device receives the globally aggregated model ${{\bf{T}}_{{\rm{FedAvg}}}}$ and updates its weights accordingly, i.e., ${\bf{T}}_{n}\leftarrow{{\bf{T}}_{{\rm{FedAvg}}}}$, $\forall n$, to synchronize all devices with a consistent model.
This iterative process continues until convergence, guided by the global objective function defined in (4). During the inference phase, each TinyML device adapts the converged model to its specific target task.
EdgeSplit is applied to a defined subset of TinyML devices. In the context of massive 6G connectivity, device heterogeneity is the expected norm, with differences in computing power, data distribution, and channel conditions being significant. In such scenarios, cluster-based PSL (CPSL) [145] offers a solution by building on the procedural structure of EdgeSplit. Instead of treating all devices as a uniform group, CPSL first clusters them according to shared characteristics (e.g., computing capabilities and channel conditions). These clusters contain devices with more manageable and stable data distributions, enabling more efficient training and improved model performance.
Figure 8: Integrated TinyML–LargeML system with EPSL.
Conventional PSL methods, such as EdgeSplit and CPSL, achieve scalability on the client side, whereas the server continues to carry a significant computational burden. EPSL [139] advances PSL by addressing this imbalance, targeting the issue of semi-scalability through a modified aggregation strategy. This strategy reduces server-side computational costs and training latency, enabling full scalability across both client and server tiers. The key steps of EPSL are illustrated in Fig. 8 and summarized below.
  • 1
    Simultaneous client-side local FP: Each TinyML device $n\in{\cal N}=\left\{{1,\ldots,N}\right\}$ independently trains on its local dataset ${{\cal D}_{n}}=\left\{{1,\ldots,{D_{n}}}\right\}$ using its local model segment, ${\bf{T}}_{n}$. The device generates smashed data ${\bf{T}}_{n}({\bf{x}}_{n})$ at the designated split layer. This step is executed in parallel across all devices.
  • 2
    FP of smashed data: Devices transmit their smashed data and corresponding labels, $\left\{{{{\bf{T}}_{n}}\left({{\bf{x}}_{n}}\right),{\bf{y}}_{n}}\right\}$, to the server. The server processes these multiple smashed data streams concurrently.
  • 3
    Server-side local FP: The server concatenates the smashed data ${\bf{S}}=\left[{{{\bf{T}}_{1}}\left({{{\bf{x}}_{1}}}\right);\ldots;{{\bf{T}}_{N}}\left({{{\bf{x}}_{N}}}\right)}\right]$, then forwards this data through its model segment $\bf{L}$ to generate the predicted output $\hat{\bf{y}}={\bf{L}}({\bf{S}})$.
  • 4
[…]
  • 5
    BP1 (aggregated gradients): The server broadcasts the aggregated activation gradients ${\bf{G}}_{a}$, computed at the split layer by performing BP on the last-layer aggregated activation gradients, to all participating TinyML devices.
  • 6
    BP2 (unaggregated activation gradients): To enable more precise updates, the server also sends the unaggregated activation gradients, denoted ${\bf{G}}_{u}$, computed at the split layer by performing BP on the last-layer aggregated activation gradients, to the respective TinyML devices.
  • 7
    Client-side local BP: Each TinyML device uses the received gradients ${\bf{G}}_{a}$ and ${\bf{G}}_{u}$ to compute its local gradient ${\bf{G}}_{n}$ and update its model segment according to ${{\bf{T}}_{n}}\leftarrow{{\bf{T}}_{n}}-{\eta_{n}}{{\bf{G}}_{n}}$.
The EPSL training process is repeated until convergence. Similarly, (4) can be used as the global objective function for the EPSL training process. During inference, each TinyML device adapts the trained model to its specific task, enabling personalized performance on its local data. Compared to standard PSL, EPSL reduces BP computation and communication complexity from ${\mathcal{O}}(N)$ to ${\mathcal{O}}(1)$, achieving full scalability [75]. EPSL also introduces a tunable aggregation ratio, denoted $r$, in the BP phase, enabling a trade-off between computational and communication efficiency while maintaining learning accuracy. When $r=0$, EPSL reverts to PSL.
Building on the PSL-based multi-level server collaboration framework introduced in Section IV-C2, we now examine parallel-server PSL and multi-hop PSL in detail. Parallel-server PSL [143] enables multiple parallel servers to process model segments concurrently for a set of resource-constrained clients. This setup introduces orchestration management and aggregation strategies that require the solution of joint optimization problems related to client-server assignments and scheduling decisions that minimize total training time. To address this NP-hard problem, we propose two scalable solutions: (i) a decomposition-based approach that exploits structural symmetry, and (ii) a lightweight alternative with low computational overhead. In contrast, multi-hop PSL [144] implements a pipelined, parallel multi-hop SL, extending the multi-hop SL approach discussed in Section IV-C1 with a parallel execution mechanism. Specifically, all participating TinyML devices simultaneously train on their local datasets using their respective model segments, generate smashed data at the split layer, and forward this data to the first server. The subsequent inference and backward paths between the servers proceed sequentially, following the same steps as in multi-hop SL. Until the BP process completes at the first server, all participating TinyML devices receive their corresponding smashed data gradients from this server, ensuring synchronization during the final aggregation phase. Unlike multi-hop SL, multi-hop PSL reduces the number of inference and backward path steps to $(2K+2)$, regardless of the value of $N$.
IV-D Federated Split Learning
Figure 9: Integrated TinyML–LargeML system with FSL.
IV-D1 Vanilla Federated Split Learning
FSL parallelizes client-side model training (see Section IV-C2) while addressing highly non-IID data across clients, combining the strengths of SL and FL [146]. FSL differs from PSL, EdgeSplit, and EPSL methods by the division of responsibilities between two cooperating servers: the main server (typically a high-performance cloud server) and the Fed server (usually an edge server running FedAvg). In FSL, the Fed server aggregates updated local models from clients using FL-based techniques. This process is illustrated in Fig. 9, including the key steps:
  1. 1
    Simultaneous client-side local FP : Each TinyML device ${\cal N}=\left\{{1,\ldots,N}\right\}$ performs local FP on its model segment, ${\bf{T}}_{n}$, using its local dataset ${{\cal D}_{n}}=\left\{{1,\ldots,{D_{n}}}\right\}$ to generate smashed data ${\bf{T}}_{n}({\bf{x}}_{n})$ at the split layer.
  2. 2
    FP of smashed data : Devices concurrently transmit the smashed data and corresponding labels $\left\{{{{\bf{T}}_{n}}\left({{\bf{x}}_{n}}\right),{\bf{y}}_{n}}\right\}$ to the main server for further centralized processing.
  3. 3
    Main-server processing: The main server processes the smashed data in both local FP and BP using either strategies of SplitFedV1 (when the server has substantial computing capacity and can efficiently execute parallel processing) or SplitFedV2 (personalized updates on the main server are essential, and client data is relatively homogeneous) [146, 147]. In SplitFedV1, each client’s smashed data is separately processed and parallelized on its model segment $\bf{L}$, where the loss value ¯$\bar{\cal L}({{{\bf{y}}_{n}},{{{\bf{\hat{y}}}}_{n}}})$ and its gradient ¯$\nabla\bar{\cal L}({{{\bf{y}}_{n}},{{{\bf{\hat{y}}}}_{n}}})$ are calculated based on the predicted output ${{\bf{\hat{y}}}_{n}}={\bf{L}}({{{\bf{T}}_{n}}({{{\bf{x}}_{n}}})})$ and ground-truth label ${{\bf{y}}_{n}}$. In the subsequent steps, the server aggregates the gradients of the smashed data and updates the server-side model using the FedAvg algorithm [106], expressed as ${\bf{L}}\leftarrow{\bf{L}}-{\eta_{S}}{{\sum\nolimits_{n\in{\cal N}}{{D_{n}}\nabla\bar{\cal L}\left({{{\bf{y}}_{n}},{{{\bf{\hat{y}}}}_{n}}}\right)}}}\Big{/}{{\sum\nolimits_{n\in{\cal N}}{{D_{n}}}}}$. In SplitFedV2, the server processes client smashed data sequentially on $\bf{L}$, following a randomly determined client order represented by the list $\cal S$ and the computation of ¯$\nabla\bar{\cal L}\left({{{\bf{y}}_{n}},{{{\bf{\hat{y}}}}_{n}}}\right)$ akin to SplitFedV1. Instead of aggregation, the server sequentially process the data in the order defined by $\cal S$ to update the global model after each local FP using the rule ${\bf{L}}\leftarrow{\bf{L}}-{\eta_{S}}\nabla\bar{\cal L}\left({{{\bf{y}}_{n}},{{{\bf{\hat{y}}}}_{n}}}\right)$ and continuing until all clients in the list $\cal S$ are processed.
  4. 4
    BP of smashed data gradients: After performing local FP and BP on the server-side model, the main server returns the evaluation of ¯$\nabla\bar{\cal L}\left({{{\bf{y}}_{n}},{{{\bf{\hat{y}}}}_{n}}}\right)$ to TinyML devices.
  5. 5
    Client-side local BP: Upon receiving the gradients from the main server, each TinyML device $n\in{\cal N}$ calculates ¯$\nabla\bar{\cal L}\left({{{\bf{T}}_{n}}\left({{{\bf{x}}_{n}}}\right)}\right)$ and performs local BP, updating its model as ${{\bf{T}}_{n}}\leftarrow{{\bf{T}}_{n}}-{\eta_{n}}\nabla\bar{\cal L}\left({{{\bf{T}}_{n}}\left({{{\bf{x}}_{n}}}\right)}\right)$.
  6. 6
    Client-side local model upload: Devices upload their updated local model to the Fed server for aggregation.
  7. 7
    Fed-server aggregation: The Fed server aggregates the client-side local models using (5) to generate a global client-side model ${\bf{T}}_{\rm{FedAvg}}$ that incorporates updates from all participating TinyML devices.
  8. 8
    Client-side global model update: ${\bf{T}}_{\rm{FedAvg}}$ is returned to all TinyML devices.
  9. 9
    Client-side local model synchronization: Device update their local model with the rule ${\bf{T}}_{n}\leftarrow{{\bf{T}}_{{\rm{FedAvg}}}}$, $\forall n$. This step ensures that all devices converge toward improved and consistent model performance.
This FSL cycle repeats iteratively until the overall model converges. During inference, each TinyML device adapts the converged model to its specific target task.
Several recent studies have explored the use of FSL for integrating TinyML and LargeML in a range of applications [148, 149, 150, 151]. For example, AdaptSFL, based on SplitFedV2, dynamically adjusts model splitting and aggregation to optimize latency and convergence on resource-constrained devices, outperforming benchmarks [148]. FedSL, also on SplitFedV2, enables robust healthcare analytics on wearables, excelling in medical imaging across diverse data [149]. Beyond standard FSL architectures,SplitGP, built on SplitFedV1, supports bidirectional training for personalized client models and generalized server models, surpassing baselines [150]. ESFL, also on SplitFedV1, enhances training stability by incorporating prior global models, using an iterative algorithm to solve its NP-hard optimization, improving efficiency over standard FL, SL, and FSL [151]. Hierarchical FSL frameworks further boost efficiency: one trains multiple edge server-client pairs with cloud aggregation for resilience [152], while another uses parallel multi-client setups for enhanced performance on constrained IoT devices [153].
IV-D2 Federated Split Foundation Models and Federated Split Fine-Tuning of Foundation Models
Federated split FMs (FSFMs) extend the FL framework by partitioning large FMs across clients and servers, enabling distributed training and inference with reduced local resource requirements [154]. Clients train or fine-tune distinct FM segments collaboratively, preserving privacy by not exposing raw data, and lowering computational overhead. Federated split fine-tuning (FedSFT) further enables client-side personalization of model components while retaining global knowledge through coordinated updates. Compared to standard FL and FFMs, FSFMs improve scalability by exchanging only compact smashed data rather than full model parameters [154]. These features position FSFMs and FedSFT as particularly suitable approaches in 6G scenarios characterized by device heterogeneity, limited computational capacity, and stringent privacy requirements.
For instance, SFPrompt is a prompt-based FedSFT framework designed for privacy-sensitive, resource-constrained edge environments [155]. By dividing the pre-trained FM between server and client and integrating soft prompts with local data pruning, SFPrompt improves FedFT efficiency and performance and significantly reduces computational and communication costs in image classification tasks. FedSplitX is a fine-grained FSFM approach that supports multiple model partition points to accommodate diverse client capabilities in heterogeneous, resource-constrained edge environments [156]. By embedding auxiliary networks at each partition, FedSplitX reduces communication overhead and latency and improves model performance for image classification tasks. FedVZ provides a defense framework against gradient inversion attacks in FSFMs using vision transformers [157]. It replaces BP with a single forward pass based on zeroth-order optimization to approximate server gradients, proving effective in heterogeneous, resource-limited environments for image classification tasks. PRINCE provides an incentive mechanism that encourages high-quality device participation from multiple FSL tenants through strategic pricing [158]. PRINCE improves FedSFT performance across a range oftasks, including image classification, sentiment analysis, speed-to-text, and question answering.
V Applications of TinyML–LargeML Integration
V-A Data Privacy and Network Security
V-A1 Toward Data Privacy
Model inversion attacks present significant privacy threats, particularly for systems trained on sensitive data. TinyML–LargeML integration enables a robust privacy-preserving architecture through on-device inference and decentralized data processing. TinyML enhances data privacy by performing local processing on edge devices [11], minimizing dependence on centralized cloud servers and reducing the exposure of raw data. In parallel, LargeML complements this architecture by aggregating and analyzing processed data without accessing the original inputs [159]. By sharing only abstracted features or model updates, a collaborative model mitigates the risks associated with model inversion attacks. TinyML also facilitates the deployment of conventional privacy-preserving techniques such as FL, allowing distributed devices to train models collaboratively without sharing raw data [11]. Differential privacy can further enhance protection by injecting noise into data or model outputs, preventing identification of individual users while maintaining meaningful analysis [160]. When combined with LargeML, more advanced FL-based techniques, such as FTL [104], FML [115], and FSL [146], enable stronger privacy protection in AI ecosystems, as detailed in Section IV.
Privacy-preserving integration is particularly crucial in healthcare applications [99, 109, 119, 149], where wearable devices embedded with TinyML algorithms can process physiological data locally. Personal health information thus remains private, while LargeML systems, hosted on secure servers, draw from aggregated insights to improve diagnostic accuracy and predictive performance without compromising raw data. Beyond healthcare, TinyML–LargeML integration benefits a range of privacy-sensitive domains. For example, in smart agriculture [101], IoT-enabled agricultural sensors equipped with TinyML analyze soil conditions, crop health, and environmental metrics on-site, protecting proprietary farming practices, while LargeML refines predictive models for climate adaptation and yield forecasting. In smart home and city [118], on-device inference captures user behavior and preferences with minimal data exposure, while LargeML enhances service personalization through cloud-based analysis.
V-A2 Toward Network Security
As reviewed in Section III-A, TinyML devices operate under stringent resource constraints, with memory and computing capacity typically two to three orders of magnitude below that of typical IoT or edge devices [161, 56, 11]. These limitations make traditional security mechanisms inadequate, and integrating TinyML and LargeML is projected to introduce a novel secure strategy. In particular, TinyML enables lightweight, real-time anomaly detection on edge devices, identifying threats with low computational cost by monitoring local activity for deviations [162]. Integrated with LargeML, which analyzes large-scale network traffic and correlates anomalies across devices [163], this forms a multi-layered defense for 6G networks. TinyML handles local threat detection, while LargeML uncovers complex attack patterns. For instance, IoTDefender [113] uses FTL across three edge servers to aggregate local models, refined by a cloud-based CNN for pre-training. During TL, models are adapted for IoT devices, enabling efficient intrusion detection despite hardware constraints.
To support efficient encryption and secure communication, TinyML enhances data security by performing lightweight encryption and decryption on edge devices, enabling secure data transmission without relying solely on cloud services [164]. However, its limited resources restrict complex encryption protocols. LargeML addresses this by managing encryption key distribution and security protocol orchestration, dynamically updating keys to strengthen communication channels [165]. This hybrid approach balances robust security with TinyML’s constraints. Fig. 10 depicts a TinyML–LargeML ecosystem enhancing 6G data privacy and network security.
Figure 10: Complementary roles of TinyML and LargeML in enhancing data privacy and network security in 6G systems.
V-B Network Management
With the rapid growth of connected IoT devices and mobile users, 5G networks have been engineered to manage high traffic volumes and support ultra-reliable low-latency communication (uRLLC), massive machine type communication (mMTC), and enhanced mobile broadband (eMBB) using technologies such as network slicing, software-defined networking (SDN), network function virtualization (NFV), and multi-access edge computing (MEC) [166]. In which, network slicing creates virtual networks on shared infrastructure, SDN and NFV enhance flexibility by decoupling control/data planes and virtualizing functions, and MEC reduces latency by processing data at the edge. As 6G emerges, user-centric, intelligent management increases complexity, rendering traditional approaches inadequate. Closed-loop automation, powered by ML and big data analytics, is now critical [167].
Against this background, TinyML–LargeML integration transforms 6G network management by combining edge-based efficiency with large-scale analytics for adaptive, intelligent systems. TinyML enables localized optimization, monitoring real-time metrics like latency and traffic to ensure low-latency and quality-of-service (QoS) in applications like smart factories [168]. LargeML processes data from multiple TinyML nodes, identifying complex patterns, predicting failures, and optimizing network-wide resources, such as rerouting traffic to maintain uRLLC, mMTC, and eMBB connectivity [169]. For instance, a Lyapunov-based adaptive split learning framework dynamically assigns split layers to TinyML devices, reducing latency and energy use [132]. Fig. 11 is an illustrative combination of TinyML’s efficiency at the edge with LargeML’s capacity for large-scale analytics, and applications of bidirectional integration across domains. In network slicing, TinyML monitors metrics for real-time resource reallocation, while LargeML predicts congestion and optimizes slice provisioning [170, 171]. In SDN, TinyML detects anomalies like DDoS attacks at the edge, and LargeML enhances traffic management by predicting congestion and optimizing routes [172, 173]. For NFV, TinyML addresses performance degradation, while LargeML automates resource scaling [174]. In MEC, TinyML reduces latency through on-device processing, complemented by LargeML’s orchestration, ensuring efficient, low-latency 6G networks [175]. For example, SplitGP [150], an FSL strategy combining personalization and generalization for training-inference optimizations. Another is SplitGP, which optimizes the TinyML device-side model for task-specific personalization and trains the MEC-side model to generalize across out-of-distribution inputs.
Figure 11: Applications of TinyML–LargeML integration in 6G network slicing, SDN, NFV, and MEC management.
V-C Intent-based Networking
Intent-based networking (IBN) is a network management paradigm that translates high-level business objectives, or human-readable “intents”, into automated network configurations [176]. Through closed-loop automation, IBN continuously verifies and enforces these intents, focusing on three core processes [177]: (i) Intent refinement, which involves the translation of user-defined intents into maintainable and adaptable network policies, (ii) Intent activation, which uses network analytics and machine learning to identify and resolve policy conflicts, and (iii) Intent assurance, which integrates with SDN controllers and orchestration systems, using telemetry to continuously monitor network functions and ensure that intents are met through ongoing adjustments.
Figure 12: IBNs with TinyML–LargeML integration.
The effectiveness of IBN depends heavily on the accurate and consistent interpretation of user inputs, such as voice, text, or other data, and the extraction of intents from these inputs. As illustrated in Fig. 12, TinyML–LargeML integration supports this goal. TinyML serves as a compact node for real-time, on-device AI processing, enabling accurate reflection of user intents and facilitating coordinated network behavior [178]. LargeML functions as a core server, aggregating intents from TinyML nodes using techniques such as TL, FTL, FFMs, SL, FSL, or FSFMs, and performing subsequent translation into actions and policies without human intervention [179]. A notable example of TinyML–LargeML synergy in IBN is an intent-based on-device AI framework [180] that facilitates collaborative AI across resource-constrained IoT devices by capturing and reflecting user intents. In this framework, TinyML manages local processing of multi-modal sensor data, for example speech-to-text conversion and an intent translation, while LargeML centrally manages policy coordination. Reducing dependence on cloud-based processing improves security, processing speed, and energy efficiency. TinyML–LargeML integration for IBN has also been applied to resource-constrained autonomous aerial vehicle swarm networks [181]. This framework combines TinyML with explainable AI to support the tamper-proof security features of blockchain and zero-trust architecture, enabling time-sensitive intent-based predictions with interpretable cognitive intelligence. It allows the detection of false communications that could distort positional data shared between aerial vehicles. The numerical results in [181] indicate a 2.34% increase in prediction accuracy, a 2.97% gain in sensitivity, latency reduction to 43.1 ms, and storage reduction to 0.0013 MB.
V-D Zero-touch Network
The zero-touch network (ZTN) is a management paradigm that uses AI, ML, and IBN to automate operations with minimal to no human input [182]. ZTN improves network efficiency by automating routine tasks such as configuration, monitoring, and troubleshooting, which lowers operational costs and frees administrators to focus on strategic planning. ZTN also improves transmission reliability by continuously analyzing traffic and user requests, adjusting network parameters in real time, resolving faults, and restoring services autonomously. ZTN also aligns network behavior with organizational policies while maintaining the flexibility to scale with increasing network demands and evolving requirements. Importantly, ZTN supports real-time threat detection and protection by integrating advanced security mechanisms.
Figure 13: ZTNs with TinyML–LargeML integration.
Achieving zero-touch automation with zero net emissions in AI-driven wireless networks requires energy-efficient TinyML–LargeML integration, low-power transceivers, lightweight AI models, and adaptive resource allocation [183]. In response, research on energy-efficient AI has accelerated, and Fig. 13 is a vision of the integrated TinyML–LargeML ecosystem for 6G ZTNs. To support autonomous solutions in self-organizing networks, TinyML optimizes neural network training on IoT devices for sustainable local learning, and LargeML uses FL and DL at base stations or clouds to reduce energy use and balance computational load [184]. In mobile networking, TinyML enables autonomous IoT node operations, and LargeML supports coverage map generation and edge AI-as-a-service (AIaaS), as seen in NanoDeploy Automator [185]. In industrial IoT, TinyML extracts data at terminal nodes, with LargeML at gateways and clouds enhancing efficiency and traffic management, exemplified by Open RAN (O-RAN)-based ZTNs for smart factories [186, 187]. Cross-layer cybersecurity frameworks like SH-CASH use AutoML to secure physical and protocol layers, with potential for TinyML integration within O-RAN architectures, ensuring service-level agreement (SLA) compliance and high performance [188].
V-E Brain-level Metaverse
The Brain-Level Metaverse (BrainMeta) integrates brain-computer interface (BCI) technology with the metaverse to create immersive virtual environments where users control avatars, manipulate objects, and navigate using brain signals [189]. Enabled by non-invasive (e.g., EEG headsets) [190] or invasive (e.g., brain implants) [191] BCI methods, BrainMeta adapts to users’ cognitive states, delivering personalized experiences for applications like virtual learning, gaming, remote work, and rehabilitation therapies for neurological conditions. It also opens employment opportunities for individuals with disabilities by supporting metaverse-based workspaces.
Figure 14: BrainMeta with TinyML and LargeML.
Despite its potential, BrainMeta remains in an early developmental stage and faces key challenges, particularly in building brain-inspired cognitive architectures constrained by size, weight, power, and processing limits [192]. BrainMeta also demands more efficient, high-performance models to accurately interpret and represent information for enhanced user interaction, cognition, content creation, and service quality in alignment with neural characteristics. These challenges can be negotiated through the integrated deployment of TinyML and LargeML, as illustrated in Fig. 14. TinyML algorithms can exploit neuromorphic computing chips [193] to support diverse message formats across computing processes. This capability enables customization of processing elements in brain-inspired communication devices, localizes memory, and increases communication bandwidth, enhancements which accelerate image recognition and classification tasks. In addition, TinyML can operate as an edge-based platform within the metaverse, using efficient TinyML processors to reduce latency and resource demands [194]. One such optimization involves winner-take-all circuits implemented with simplified leaky-integrate-and-fire neurons on field-programmable gate arrays. LargeML-based approaches [9], such as cloud–edge–end collaborative models, mobility-aware pre-rendering, and diffusion model-based adaptive rendering, can be improve data collection from TinyML devices and enhance rendering tasks in the metaverse. The process involves extraction of generalized representations from extensive EEG data, segmentation of data into channel arrays, and application of vector-quantized neural spectrum prediction to train and update neural analyzers [195]. AMFL, an adaptive, resource-efficient FL framework designed for metaverse environments, offers another example [196]. By combining FL and DRL strategies within BrainMeta, AMFL mitigates the adverse effects of non-IID data, reduces resource demands, and improves the QoE for human-centric, resource-constrained augmented reality devices.
VI Challenges and Future Research Directions
VI-A Standardization
5G faces challenges with dynamic user demands and traffic, while 6G aims to enable sustainable applications like holography and the metaverse through TinyML–LargeML integration [197]. This integration is designed for user-centric services that adapt to changing behaviors and networks. However, standardization issues impede the development of a fully integrated ecosystem. The ITU-R identifies RAN slicing as crucial for 6G virtual network sharing [198], requiring a unified O-RAN architecture that supports diverse air interfaces and AI functions. This architecture should offer customizable service parameters, but shared infrastructure raises concerns about transparency, reliability, and privacy. Standardization bodies are advancing AI-driven wireless architectures. 3GPP’s Releases 17 [199] and 18 [200] enhance 5G RAN with AI for intelligence and efficiency, while the O-RAN Alliance integrates AI for orchestration and control [201, 202].
However, challenges remain in integrating TinyML and LargeML across the network lifecycle. Proposals include a secure TinyML edge AI architecture [60], explainable AI for decision transparency [203], and a testbed for trustworthy network management [204]. Sustainability efforts are supported by an energy cost-of-AI metric [205]. At a higher level, a shift to a multi-tier mNode model is proposed [197], featuring distributed nodes and network AI logic. Additionally, a digital twin framework for AI-RAN can manage digital twins and enable real-time adaptability in dense networks [206].
VI-B Resource Management and Orchestration
6G networks face resource management and orchestration challenges, including wireless channel uncertainty, ISAC device proliferation, and carbon emission reduction. These drive AI/ML-based frameworks to optimize spectrum, power, and time allocation, enhancing interference management, energy efficiency, and reliability over traditional methods like channel estimation and beamforming [207]. However, integrating AI/ML introduces data privacy risks, computational overhead, and scalability issues. Adaptive algorithms like DRL with diffusion models [207] and distributed learning like FTL with blockchain incentives [208] can mitigate these by preserving privacy and reducing latency.
Deploying DL with TinyML faces obstacles [209]: (i) computational trade-offs, (ii) vanishing gradients needing hardware acceleration, (iii) memory constraints for semantic communication, (iv) reliance on scarce real-world data, and (v) dynamic environmental variability. LargeML, using distributed neural networks and collaborative learning, encounters issues like: (i) data segmentation for collective learning, (ii) transmission errors in FL affecting accuracy, and (iii) stringent privacy and coordination demands. Future research into model-agnostic meta-learning [210] or accretionary learning [211] could address these challenges.
At a broader level of cross-domain interaction and orchestration, AI will be integrated into 6G network slicing, leveraging NFV for resource virtualization and SDN for centralized management. A central SDN controller handles planning and provisioning, while AI-enhanced local controllers manage operations, ensuring cost-effective, QoS-compliant network slices. However, cross-domain AI coordination poses challenges, including those outlined in [212] as follows: (i) Data/knowledge coordination, (ii) model training coordination, (iii) inference and decision coordination, and (v) computing resource coordination.
Beyond technical integration, managing and orchestrating TinyML and LargeML in 6G network slices is challenging due to the shift from 5G’s centralized learning to distributed paradigms like FL, swarm learning, multi-intelligent RL, TL, meta-learning, and SL. The choice of learning approach depends on node capabilities, dataset traits, computing resources, model complexity, latency, and network conditions. Supporting diverse distributed methods increases orchestration complexity. To address this, AI nodes should adopt specialized roles, such as horizontal FL for decentralized data, hierarchical FL for reduced communication, TL for efficient training, vertical FL for cross-domain learning, SL for uneven resources, multi-intelligent RL for localized collaboration, and swarm learning for secure joint training.
VI-C Advanced Security and Privacy-Preserving Techniques
As TinyML–LargeML integration becomes central to next-generation wireless network developments, the security and privacy landscape will continue to grow in complexity. Key concerns include adversarial attacks, data security, model trustworthiness and integrity, and user privacy.
TinyML faces security challenges due to limited edge device resources, making devices like traffic sensors, smartphones in FL, and wearables vulnerable to adversarial attacks. Adversarial training [213] enhances TinyML’s resilience to malicious inputs, while homomorphic encryption [214] ensures data confidentiality and integrity. Privacy risks from sensitive edge data can be mitigated using differential privacy, which adds noise to obscure information [215], and advanced encryption to secure model parameters during transmission [216].
LargeML deployment also requires robust privacy and security due to its broader scope and multi-stakeholder involvement. Model distillation [33] transfers LargeML knowledge to TinyML, reducing computational needs while maintaining performance. Integrating LargeML with GANs [217] simulates cyberattacks to improve resilience. Blockchain ensures model integrity by hashing and recording versions, maintaining TinyML model authenticity [215]. Advanced privacy-preserving techniques further secure LargeML systems: (i) privacy-preserving data shuffling, (ii) differential private synthetic data generation, iii) private aggregation of teacher ensembles, (iv) Rényi differential privacy for generative modeling, (v) cross-Silo FL with secure data sharding, and (vi) decentralized privacy-preserving knowledge distillation.
VI-D Real-time, Lightweight Intelligence
Beyond 5G and 6G networks aim to automate M2M operations by transforming conventional communication systems into AI-centric architectures, ultimately transitioning toward human-centric networks through advanced AI integration. To meet the growing demands for improved user experience and reduced carbon emissions, the integration of TinyML and LargeML into AI-enabled systems creates new opportunities for real-time, lightweight intelligence yet also introduces distinct technical challenges.
TinyML applications vary in communication protocols, latency requirements [11], and device specifications. For example, Arduino Nano (64 MHz, 1 MB flash memory, 256 KB RAM) and Raspberry Pi (up to 1.8 GHz, 32 GB flash memory, 8 GB RAM) reflect a range of processing capabilities [4]. This variability highlights the need for compact, unified TinyML frameworks that support flexible plug-and-play deployment, provide efficient processing, and maintain lightweight runtimes [218]. In contrast to TinyML, LargeML typically processes vast input volumes from multiple domains, each requiring tailored prompt engineering and model-splitting approaches. To deliver real-time responses with accurate and reliable outputs, LargeML often depends on extensive hyperparameter tuning and multi-step processing. These operations result in prolonged training times, high power consumption, and significant memory requirements [16]. Consequently, future LargeML should adopt adaptive prompt engineering strategies, such as model-adaptive prompts [219] and lightweight tuning paradigms for pluggable prompting [220], to improve explainability and accelerate processing and inference without disruption to the life cycle pipeline.
However, the limited availability of unified platforms and tools for managing workflows, coordinating input data, and ensuring target environment compatibility poses additional challenges in TinyML–LargeML integration. For example, LLMs require specialized ML libraries, compilers, and runtimes for code generation, text processing, and prompt execution tasks, each with varying input formats, volumes, and execution times [25]. Iterative error-handling and refinement mechanisms are often required to improve output reliability. Thus, in addition to concentrating on optimizing TinyML and LargeML design, comprehensive guidelines and evaluation frameworks are essential to define task-specific roles and system requirements for each deployment and operation phase [221].
VI-E AI-Native and Collective AI for 6G
The concept of AI-Native redefines system design by positioning AI at the core of operations rather than treating it as supplementary feature, as seen in autonomous vehicles where AI manages navigation, obstacle detection, and collision avoidance. As defined in [222], future AI-native systems aim to unify five critical capabilities: solving complex problems via data-driven AI for seamless operations, adapting to dynamic environments, combining foundational knowledge with new insights for continuous learning, ensuring trustworthy and transparent AI functions, and enabling resilient, cognitive autonomous networks. Nevertheless, integrating TinyML and LargeML into AI-Native systems poses challenges: (i) coordinating resource-constrained TinyML with computationally robust LargeML for zero-touch operations, (ii) defining their roles across heterogeneous infrastructures to have a unified framework, (iii) ensuring interoperable learning to support continuous learning across the network, (iv) managing AI life cycles for reliability and fairness, and (v) optimizing joint performance while maintaining TinyML accuracy.
Collective AI involves multi-agent collaboration, where agents contribute to shared goals through local training without direct communication [223]. At the platform level, collective reinforcement learning enables agents to share models, thereby achieving short training time, redundant learning avoidance, and robust models from limited samples [224]. At a broader implementation level, collective AI swarm intelligence supports IoT applications in static and dynamic environments [225], but it requires designing a unified framework for collective dynamics and reasoning across heterogeneous systems.
As the integration of TinyML and LargeML progresses within collective AI systems, new opportunities will emerge alongside specific challenges: (i) Crowd retention: Varying agent contributions require robust incentive mechanisms to sustain participation; (ii) Spectrum utilization: Dense IoT networks strain spectrum resources, necessitating optimized bandwidth allocation, especially with 6G; (iii) Security, trust, and privacy: Sharing training outputs raises concerns, with blockchain’s resource demands limiting its use on constrained devices, while human-AI collaboration needs better social cue interpretation, requiring advanced cognitive architectures; and (iv) Ethical issues: Balancing TinyML and LargeML benefits with risks involves addressing transparency, bias, and regulatory gaps. Errors in LargeML outputs could have significant consequences, underscoring the need for legal frameworks to ensure accountability and trust.
VII Conclusion
TinyML–LargeML integration is set to revolutionize the development of advanced AI solutions in 6G and future networks. These technologies address the critical issues of resource allocation, network management, security, and privacy. This survey provides a comprehensive examination of the convergence of TinyML and LargeML technologies, beginning with a detailed review of design considerations, operational mechanisms, and performance evaluation methods. The survey then investigates the motivations and technical requirements driving the convergence of these systems toward 6G and beyond, presenting efficient bidirectional integration strategies and highlighting application scenarios. This extensive review identifies the challenges persistent throughout the literature and outlines numerous promising research directions, including standardization, resource orchestration, advanced security and privacy-preserving techniques, real-time and lightweight AI processing, and the adoption of AI-native and collective AI paradigms. Sustained research in this domain is essential. This survey aims to support and inform the ongoing work in providing a foundation for future innovations and advancements in 6G systems.